### v1

# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

Abstract
In Deep Reinforcement Learning models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence. Additionally, due to the non-stationarity of the objective function, the best-performing learning rate can change over the training steps. To adapt the learning rate, a standard technique consists of using decay schedulers. However, these schedulers assume that the model is progressively approaching convergence, which may not always be true, leading to delayed or premature adjustments. In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that selects the learning rate based on the agent’s performance during training. LRRL is based on a multi-armed bandit algorithm, where each arm represents a different learning rate, and the bandit feedback is provided by the cumulative returns of the RL policy to update the arms’ probability distribution. Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms.
1 Introduction
Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.33, 19, 39, 40). While ANNs were previously used as value function approximators [29], the introduction of Deep Q-Networks (DQN) by [24, 25] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
The experience replay (see 22) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation. The target network further stabilizes the learning process by periodically copying the parameters of the learning network. This strategy is crucial because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence. By leveraging the target network, gradient steps are directed towards a periodically fixed target, ensuring more stability in the learning process. Additionally, the learning rate hyperparameter controls the magnitude of these gradient steps in optimizers such as the stochastic gradient descent algorithm, affecting the training convergence.
The learning rate is one of the most important hyperparameters, with previous work demonstrating that decreasing its value during policy finetuning can enhance performance by up to 25% in vanilla DQN [3]. Determining the appropriate learning rate [^1] The terms “learning rate” and “step-size” are often used interchangeably in the literature and they technically refer to the same concept. is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 14, 7, 42). However, finding a learning rate value that improves the model performance requires extensive and computationally expensive testing. In order to adapt its initial choice during training, optimizers such as Adam [17] and RMSProp [37] employ an internal scheme that dynamically adjusts the learning rate, considering, for instance, past gradient information. Nevertheless, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
Standard learning rate schedulers typically decrease the learning rate based on training progress using, e.g., linear or exponential decay strategies [31, 42]. In the context of RL, this approach can lead to premature or delayed learning rate adjustments, which may hinder the agent’s ability to learn. Unlike supervised learning, RL usually involves generating data by trading off between exploration (discovery of new states) and exploitation (refining of the agent’s knowledge). As the policy improves, the data distribution encountered by the agent becomes more concentrated, but this evolution occurs at a different pace than the overall training progress. For instance, some environments require extensive exploration due to the sparseness of rewards, while others need more exploitation to refine the policy to the complexity of the task. Consequently, a more sophisticated decaying learning rate strategy that accounts for policy performance rather than training steps can significantly enhance learning in deep RL.
In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a method to select the learning rate on the fly for deep RL. Our approach acknowledges that different learning phases require different learning rates, and as such, instead of scheduling the learning rate decay using some blanket approach, we dynamically choose the learning rate using a Multi-Armed Bandit (MAB) algorithm, which accounts for the current policy’s performance. Our method has the advantage of being algorithm-agnostic and applicable to any optimizer, although the results show that it works best when coupled with Adam. We conduct experiments on our approach using baselines provided in the Dopamine framework (see 9). Our results focus on exploiting different settings for LRRL to illustrate its robustness under many possible configurations. Our main contributions are the following:
  • We introduce LRRL, the first approach, to our knowledge, that leverages a multi-armed bandit algorithm to select the learning rate dynamically in deep RL. Our results demonstrate that LRRL achieves competitive performance with or superior to standard deep RL algorithms using fixed baselines or traditional learning rate schedulers.
  • Our results show that LRRL significantly reduces the need for hyperparameter optimization by dynamically selecting from a set of possible learning rates using a multi-armed bandit approach. This method mitigates the need for exhaustive techniques like grid search, as it efficiently adapts the learning rate during training in a single run, rather than requiring multiple runs to test each learning rate individually.
  • We assess the robustness of our method by employing the Adam and RMSProp optimizers with different sets of arms. We also compare the results using stochastic and adversarial multi-armed bandit algorithms in Appendix A.2.
2 Related Work
Multi-armed bandit for (hyper)parameter selection.
Deep RL is known to be overly optimistic in the face of uncertainty [27], and many works have addressed this issue by proposing conservative policy updates [38, 13, 2]. However, when the agent is able to interact with the environment, this optimism can encourage exploration, potentially leading to the discovery of higher returns. Building on this idea, [26] use an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process. In order to select the learning rate for stochastic gradient MCMC, [10] employ an algorithm based on Successive Halving [16, 15], a MAB strategy that promotes promising arms and prunes suboptimal ones over time. In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [21, 32]. A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space. Close to our work, [23] propose Adam with Bandit Sampling (ADAMBS), which employs the Exponential-weight algorithm for Exploration and Exploitation [6] to enhance sample efficiency by incorporating importance sampling within the Adam optimizer. While ADAMBS prioritizes informative samples, our method focuses on RL tasks by dynamically adjusting the learning rate, accelerating learning when the policy is far from optimal, and slowing it down as it converges.
Learning rate adapters/schedulers.
Optimizers such as RMSProp [37] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [17] also incorporates momentum to smooth gradient steps. However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see  12). Increment-Delta-Bar-Delta (IDBD), introduced by  [34], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [43]. Learning rate schedulers with time decay [31, 42] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning. A meta-gradient reinforcement learning is proposed in [41], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function. Our work differs from these methods by employing a multi-armed bandit approach to dynamically select the learning rate over the training process, specifically targeting RL settings.
3 Preliminaries
This section introduces the Reinforcement Learning and Multi-Arm Bandits frameworks, defining supporting notation.
3.1 Deep Reinforcement Learning
An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.35 for details). In RL, starting from an initial state $s_{0}$, a learner called agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$. In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(s_{t},a_{t},\,\cdot\,)$. The agent’s objective is to learn a policy $\pi:\mathcal{S\times A}\rightarrow[0,1]$ which maps a distribution of actions given the current state, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr{]}$.
To learn how to perform a task, value function-based algorithms coupled with ANNs [24, 25] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$. By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function $\mathcal{J}(\theta)$ defined by:
$\mathcal{J}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\Bigl{[}r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})-Q_{\theta}(s,a)\Bigr{]}^{2}\,,$ (1)
where $\theta^{-}$ are the target network parameters used to compute the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$. The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$. The minimization, and hence, the update of the parameters $\theta$, is done according to the optimizer’s routine. A simple possibility is to use stochastic gradient descent using mini-batch approximations of the loss gradient:
$\theta_{n+1}\leftarrow\theta_{n}-\eta\nabla_{\theta}\mathcal{J}(\theta_{n})\,,\quad\text{where}\quad\nabla_{\theta}\mathcal{J}(\theta_{n})\approx\frac{1}{|\mathcal{B}|}\sum_{(s,a,r,s^{\prime})\in\mathcal{B}}2\bigl{(}Q_{\theta}(s,a)-y\bigr{)}\nabla_{\theta}Q_{\theta}(s,a)$ (2)
with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called the learning rate. Unlike in supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time. This continuous evolution introduces instability during the learning process, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$. However, stability also depends on how the parameters $\theta$ change during each update. This work aims to control these changes by dynamically selecting the learning rate $\eta$ over the training steps.
3.2 Multi-Armed Bandit
Multi-Armed Bandits (MAB) provide an elegant framework for making sequential decisions under uncertainty (see for instance 20). MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$. Like RL, MAB algorithms must balance the trade-off between exploring arms that have been tried less frequently and exploiting arms that have yielded higher rewards up to time $n$.
To account for the non-stationarity of the RL rewards, we will consider in this work the MAB setting of adversarial bandits [6]. In this setting, at each round $n$, the agent selects an arm $k_{n}$ according to some distribution $p_{n}$ while the environment (the adversary) arbitrarily (e.g., without stationary constraints) determines the rewards $f_{n}(k)$ for all arms $k\in\mathcal{K}$. MAB algorithms are designed to minimize the pseudo-regret $G_{N}$ after $N$ rounds defined by:
$G_{N}=\max_{k\in\{1,\dots,K\}}\,\mathbb{E}\Biggl{[}\sum_{n=1}^{N}f_{n}(k)-\sum_{n=1}^{N}f_{n}(k_{n})\Biggr{]}\,,$
where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}f_{n}(k)}$ represents the accumulated reward of the single best arm in hindsight, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm. A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$: this guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
4 Dynamic Learning Rate for Deep RL
In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic Learning Rate for deep Reinforcement Learning (LRRL). LRRL is a meta-learning approach designed to dynamically select the learning rate in response to the agent’s performance. LRRL couples with stochastic gradient descent optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm. As the agent interacts with the environment, the average of observed rewards is used as bandit feedback to guide the selection of the most appropriate learning rate throughout the training process.
4.1 Selecting the Learning Rate Dynamically
Our problem can be framed as selecting a learning rate $\eta$ for policy updates —specifically, when updating the parameters $\theta$ after $\lambda$ interactions with the environment.
Before training, a set $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ of $K$ learning rates are defined by the user. Then, during training, a MAB algorithm selects, at every round $n$ —that is, at every $\kappa$ interactions with the environment— an arm $k_{n}\in\{1,\dots,K\}$ according to a probability distribution $p_{n}$ defined based on previous rewards, as explained in the next section. The parameters $\theta$ are then updated using the sampled learning rate $\eta_{k_{n}}$. The steps involved in this meta-learning approach are summarized in Algorithm 1.
Algorithm 1 dynamic Learning Rate for deep Reinforcement Learning (LRRL)
Parameters:
    Set of learning rates $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$
    Number of episodes $M$
    Horizon length $T$
    Update window $\lambda$ for the learning network $\theta$
    Update window $\tau$ for the target network $\theta^{-}$
    Update window $\kappa$ for arm probabilities $p$
Initialize:
    Parameters $\theta$ and $\theta^{-}$
    Arm probabilities $p_{0}\leftarrow(\tfrac{1}{K},\dots,\tfrac{1}{K})$
    MAB round $n\leftarrow 0$
    Cumulative reward $R\leftarrow 0$
    Environment interactions counter $C\leftarrow 0$
for episode $m=1,2,3,\ldots,M$ do
     for timestep $t=1,2,3,\ldots,T$ do
         Choose action $a_{t}$ following the policy $\pi_{\theta}(s)$ with probability $1-\epsilon$triangleright $(\epsilon$-greedy strategy)
         Play action $a_{t}$ and observe reward $r_{t}$
         Add $r_{t}$ to cumulative reward $R\leftarrow R+r_{t}$
         Increase environment interactions counter $C\leftarrow C+1$
         if $C\mod\lambda\equiv 0$ then
              if $C\geq\kappa$ then
                  Compute average of the last $C$ rewards $f_{n}\leftarrow\frac{R}{C}$
                  Increase MAB round $n\leftarrow n+1$
                  Compute weights $w_{n}$ and arm probabilities $p_{n}$ using Equations (3, 4)
                  Sample arm $k_{n}$ with distribution $p_{n}$
                  Reset $R\leftarrow 0$ and $C\leftarrow 0$
              end if
              Update network parameter $\theta$ using the optimizer update rule with learning rate $\eta_{k_{n}}$
              Every $\tau$ steps update the target network $\theta^{-}\leftarrow\theta$
         end if
     end for
end for
Note that the same algorithm might be used with learning rates schedulers, that is with $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ where $\eta_{k}:\mathbb{N}\to\mathbb{R}_{+}$ is a predefined function, usually converging towards 00 at infinity. If so, the learning rate used at round $n$ of the optimization is $\eta_{k_{n}}(n)$.
4.2 Updating the Probability Distribution
As we expect that the agent’s performance —and hence, the cumulative rewards— will improve over time, the MAB algorithm should receive non-stationary feedback. To take this non-stationary nature of the learning into account, we employ the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction). At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively. Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to improvements in policy performance.
Specifically, after picking arm $k_{n}$ at round $n$, the RL agent interacts $C$ times with the environment and the MAB algorithm receives a feedback $f_{n}$ corresponding to the average reward of those $C$ interactions. Based on [26], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
$f_{n}^{\prime}=f_{n}-\frac{1}{j}\sum_{i=0}^{j-1}f_{n-i}\,.$
The improvement in performance allows computation of the next weights $w_{n+1}$ as follows, where initially $w_{1}=(0,\ldots,0)$:
$\forall k\in\{1,\ldots,K\},\quad w_{n+1}(k)=\begin{cases}\delta\,w_{n}(k)+\alpha\,\frac{f^{\prime}_{n}}{e^{w_{n}(k)}}&\text{if }k=k_{n}\\ \delta\,w_{n}(k)&\text{otherwise}\,,\end{cases}$ (3)
where $\alpha>0$ is a step-size parameter. The distribution $p_{n+1}$, used to draw the next arm $k_{n+1}$, is
$\forall k\in\{1,\ldots,K\},\quad p_{n+1}(k)=\frac{e^{w_{n+1}(k)}}{\sum_{k^{\prime}=1}^{K}e^{w_{n+1}(k^{\prime})}}\,.$ (4)
This update rule ensures that as the policy $\pi_{\theta}$ improves, the MAB algorithm continues to favor learning rates that are most beneficial under the current policy performance, thereby effectively handling the non-stationarity inherent in the learning process.
5 Experiments
In the following sections, we investigate whether combining LRRL with Adam or RMSProp —two widely used optimizers— can improve cumulative returns in deep RL algorithms. To assess this, we compare LRRL against learning methods with and without schedulers using the baseline implementation of DQN provided in Dopamine [9]. We test LRRL under different configurations and Atari games, reporting the average and standard deviation of returns over 5 runs. Details on the evaluation metrics and hyperparameters used in these experiments are summarized in Appendix B.
5.1 Comparing LRRL with Standard Learning
In our first experiment, we consider a set of $5$ learning rates, and compare the performance of 5 configurations of LRRL, each of them using a subset of those learning rates, against the DQN algorithm reaching best performance, in terms of maximum average return, among the $5$ possible learning rates choices (see Figure 5 in Appendix A.1). More precisely, the set of learning rates is
$\mathcal{K}(5)=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}\,.$
The whole set $\mathcal{K}(5)$ is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) and 3 taking the lowest/middle/highest ($\mathcal{K}_{\text{sparse}}(3)$) learning rate values. All experiments use the Adam optimizer, and we report the return based on the same number of environment iterations.
Results are gathered in Table 1 and illustrated in Figure 1 along with four Atari games. They show that LRRL outperforms standard learning in two out of four tasks while remaining competitive in the others. Notably, LRRL not only matches or exceeds performance but also reduces the need for extensive parameter tuning by incorporating multiple learning rates in a single run. However, as the results indicate, the $3$-arm bandit variants exhibit different behaviors during the learning process, suggesting that the choice of the number of arms and their values still requires task-specific tuning to achieve good performance.
Figure 1: A comparison between 5 configurations of LRRL with various learning rate subsets and the DQN algorithm reaching the best performance among possible learning rates.
Game DQN LRRL $\mathcal{K}_{\text{lowest}}(3)$ LRRL $\mathcal{K}_{\text{middle}}(3)$ LRRL $\mathcal{K}_{\text{highest}}(3)$ LRRL $\mathcal{K}_{\text{sparse}}(3)$ LRRL $\mathcal{K}(5)$
Asteroids 1 085 $\pm$ 63 1 065 $\pm$ 70 1 079 $\pm$ 48 1 061 $\pm$ 59 1007 $\pm$ 127 1 085 $\pm$ 88
Breakout 217 $\pm$ 14 220 $\pm$ 24 201 $\pm$ 6 177 $\pm$ 8 270 $\pm$ 13 253 $\pm$ 20
Pong 19 $\pm$ 0.4 18 $\pm$ 0.3 19 $\pm$ 0.9 19 $\pm$ 0.5 19 $\pm$ 0.5 19 $\pm$ 0.8
Seaquest 5 881 $\pm$ 1 533 5 905 $\pm$ 2 557 6 135 $\pm$ 2 229 6 231 $\pm$ 1 802 8 920 $\pm$ 2 759 6 799 $\pm$ 2 060
Table 1: Max average return (best in bold if significantly better than others) and its standard deviation for 4 Atari games.
To illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves, Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$. In most of the tested environments, LRRL behaves similarly to time-decay schedulers by selecting higher learning rates during the early stages of training, gradually shifting toward arms with lower rates as training progresses. The exception is in the Pong environment, where the model converges after only a few iterations, resulting in a more uniform probability distribution across the set of arms.
Figure 2: Systematic sampling of normalized learning rates and returns over the training steps using LRRL $\mathcal{K}(5)$ with Adam optimizer, through a single run. For each episode, we show the selected learning rate using different colors.
5.2 Combining and Comparing Schedulers with LRRL
Next, we consider using LRRL combined with learning rate schedulers. Specifically, we employ schedulers with exponential decay rate of the form $\eta(n)=\eta_{0}\times e^{-d\,n}$, where $\eta_{0}$ is a fixed initial value (common to each scheduler and equal to $6.25\times 10^{-5}$ in our experiment), $d$ is the exponential decay rate and $n$ is the number of policy updates (i.e., of MAB rounds). We define a set of $3$ schedulers $\mathcal{K}_{s}$, where each arm represents a scheduler using a different decay rate $d=\{1,2,3\}\times 10^{-7}$, and compare the results of LRRL with each scheduler individually, using the Adam optimizer.
Figure 3 and Table 2 show that LRRL combined with schedulers can substantially increase final performance compared to using exponential decay schedulers for some environments while remaining competitive for others. The dashed black line represents the max average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by [4], who linearly decay the learning rate to 00.
Figure 3: A comparison between LRRL with arms as schedulers and schedulers individually.
Game DQN $d=1\times 10^{-7}$ $d=2\times 10^{-7}$ $d=3\times 10^{-7}$ LRRL ($\mathcal{K}_{s}$)
Asteroids 1 028 $\pm$ 53 1 025 $\pm$ 52 1 013 $\pm$ 56 1 063 $\pm$ 82 1 015 $\pm$ 33
Breakout 144 $\pm$ 12 149 $\pm$ 11 151 $\pm$ 7 144 $\pm$ 12 233 $\pm$ 19
Seaquest 5 881 $\pm$ 1 533 5 284 $\pm$ 1 134 5 793$\pm$ 1 225 6 612 $\pm$ 835 13 864 $\pm$ 3 581
Video Pinball 410 186 $\pm$ 193 328 388 768 $\pm$ 195 150 450 015 $\pm$ 178 876 362 645 $\pm$ 172 169 388 308 $\pm$ 103 862
Table 2: Max average return (best in bold) and its standard deviation for 4 Atari games.
5.3 RMSProp Optimizer and more Environments
Another widely used optimizer for training deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism. Adam builds upon RMSProp by retaining exponential moving averages to give more weight to recent gradients while incorporating momentum. Although the standard RMSProp does not feature momentum, we found that adding momentum to RMSProp can increase both the performance of DQN and LRRL, aligning with findings in the literature [28, 4].
In the following experiment, we compare the performance of RMSProp with Nesterov’s momentum (RMSProp-M), and Adam when coupled with either LRRL or the best-performing single learning rate when using DQN. As shown in Figure 4 and Table 3, LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN. Moreover, LRRL (RMSProp-M) underperforms compared to DQN without LRRL in two out of three tasks due to its slow convergence despite better jumpstart performance. Future work should investigate whether this slow convergence is linked to factors such as the environment’s stochasticity or the optimizer’s features such as the absence of bias correction in the first and second moment estimates.
Figure 4: A comparison between Adam and RMSProp with momentum, using either DQN or LRRL.
Game DQN (RMSProp-M) LRRL (RMSProp-M) DQN (Adam) LRRL (Adam)
Asterix 11 464 $\pm$ 2 848 6 499 $\pm$ 927 12 561 $\pm$ 1 245 15 017 $\pm$ 3 892
Ms. Pacman 3 301 $\pm$ 310 2 696 $\pm$ 248 3 232 $\pm$ 114 3 310 $\pm$ 231
Space Invaders 1 490 $\pm$ 132 2 712 $\pm$ 73 2 874 $\pm$ 319 3 641 $\pm$ 1 030
Table 3: Max average return (best in bold) and its standard deviation for 3 Atari games.
6 Conclusion
In this work, we introduced dynamic Learning Rate for Deep Reinforcement Learning (LRRL), a meta-learning approach for selecting the optimizer’s learning rate on the fly. We demonstrated empirically that combining LRRL with the Adam optimizer could significantly enhance the performance of the value-based algorithm DQN, outperforming baselines and learning rate schedulers in some tasks while remaining competitive in others. Furthermore, by employing a multi-armed bandit algorithm, LRRL reduces the need for extensive hyperparameter tuning, as it explores a set of learning rates in a single run with minimal extra computational overhead.
While this work focused on dynamically selecting the best-performing learning rate, future investigations could extend LRRL ideas to other critical hyperparameters, such as mini-batch size, which also plays a key role in the model’s convergence. Moreover, although LRRL selects learning rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [30, 18, 1].
Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation 2024-A0171015651 made by GENCI. HD and FF acknowledge funding from Inria Project WOMBAT. LFS is supported by a Discovery Early Career Researcher Award from the Australian Research Council (DE240101190)
Appendix A Supplementary Experiments
A.1 Baseline Evaluation with Varying Learning Rates
To establish a baseline to compare learning without our approach LRRL, we run individual arm values as baseline learning rate using the Adam optimizer. The results presented in Figure 5 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to the worst jumpstart performance and slow convergence.
Figure 5: DQN performance using Adam optimizer with varying learning rates across 4 Atari games.
A.2 A Comparison between Multi-Armed Bandits Algorithms
Adversarial MAB algorithms are designed for environments where the reward distribution changes over time. In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown probability distributions. To validate our choice and address our method’s robustness, we compare Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [5]. MOSS trade-off exploration-exploitation by pulling the arm $k$ with highest upper confidence bound given by:
$B_{k}(n)=\hat{\mu}_{k}(n)+\rho\,\sqrt{\frac{\max\Bigl{(}\log{\frac{n}{Kn_{k}(n)}},0\Bigr{)}}{n_{k}(n)}}$
where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to timestep $t$.
In Figure 6, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balance exploration and exploitation in MOSS. Additionally, the bandit feedback used in MOSS is the average cumulative reward $\frac{R}{C}$. The results indicate that while Exp3 performs better overall, MOSS can still achieve competitive results depending on the amount of exploration, demonstrating the robustness of our approach regarding the MAB algorithm employed.
Figure 6: A comparison between adversarial and stochastic MAB algorithms.
Appendix B Experiments Details
B.1 Evaluation Terminology
In this section, we describe the evaluation metrics that can be used to evaluate agent’s performance as it interacts with an environment. Based on Dopamine, we use the evaluation step-size “iterations”, which is defined as a predetermined number of episodes. Figure 7 illustrates the evaluation metrics used in this work, as defined in [36]:
  • Max average return: The highest average return obtained by an algorithm throughout the learning process. It is calculated by averaging the outcomes across multiple individual runs.
  • Final performance: The performance of an algorithm after a predefined number of interactions. While two algorithms may reach the same final performance, they might require different amounts of data to do so. This metric captures the efficiency of an algorithm in reaching a certain level of performance within a limited number of interactions. In Figure 7, the final performance overlaps with the max average return, represented by the black dashed line.
  • Jumpstart performance: The performance at the initial stages of training, starting from a policy with randomized parameters $\theta$. In Figure 7, Algorithm B exhibits better jumpstart performance but ultimately achieves lower final performance than Algorithm A. A lower jumpstart performance can result from factors such as a lower learning rate, although this work demonstrates that this does not necessarily lead to worse final performance.
Figure 7: Performance curves of two RL algorithms (adapted from taylor2009transfer [36]).
B.2 Further Experimental Details
In the following, we list the set of arms, optimizer and the bandit step-size used in each experiment.
Section 5.1 – Comparing LRRL with Standard Learning.
  • Optimizer: Adam
  • Bandit step-size: $\alpha=0.2$
  • Considered sets of learning rates:
    $\mathcal{K}(5)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
    $\mathcal{K}_{\text{lowest}}(3)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
    $\hskip 25.03856pt\mathcal{K}_{\text{middle}}(3)=\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
    $\hskip 84.50471pt\mathcal{K}_{\text{highest}}(3)=\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
[…]
Section 5.2 – Combining and Comparing Schedulers with LRRL.
  • Optimizer: Adam
  • Bandit step-size: $\alpha=0.2$
  • Initial learning rate of schedulers: $\eta_{0}=6.25\times 10^{-5}$
Section 5.3 – RMSProp Optimizer and More Environments.
  • Optimizer: RMSProp for baselines
  • Bandit step-size: $\alpha=0.2$
  • LRRL set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
Section A.2 – A Comparison between Multi-Armed Bandit Algorithms.
  • Optimizer: Adam
  • MOSS/Exp3 set of learning rates: $\mathcal{K}=\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
B.3 Hyperparameters
In this section, we outline the hyperparameters used in the experiments. The optimizers from the Optax library [11] are employed alongside the JAX [8] implementation of the DQN algorithm [24, 25], as provided by the Dopamine framework [9].
Hyperparameter Setting
Sticky actions True
Sticky actions probability 0.25
Discount factor ($\gamma$) 0.99
Frames stacked 4
Mini-batch size ($\mathcal{B}$) 32
Replay memory start size 20000
Learning network update rate ($\lambda$) 4 steps
Minimum environment steps ($\kappa$) 1 episode
Target network update rate ($\tau$) 8000 steps
Initial exploration ($\epsilon$) 1
Exploration decay rate 0.01
Exploration decay period 250000 steps
Environment steps per iteration 250000 steps
Reward clipping [-1, 1]
Network neurons per layer 32, 64, 64
Hardware V100 GPU
Adam hyperparameters
$\beta_{1}$ decay 0.9
$\beta_{2}$ decay 0.999
Eps 1.5e-4
RMSProp hyperparameters
Decay 0.9
Momentum (if True) 0.999
Centered False
Eps 1.5e-4
Table 4: Hyperparameters

### v2

@@
 Abstract
-In Deep Reinforcement Learning models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence.
+In deep Reinforcement Learning (RL) models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence.
 Additionally, due to the non-stationarity of the objective function, the best-performing learning rate can change over the training steps.
@@
 LRRL is based on a multi-armed bandit algorithm, where each arm represents a different learning rate, and the bandit feedback is provided by the cumulative returns of the RL policy to update the arms’ probability distribution.
-Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms.
+Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms for some tasks.
 1 Introduction
-Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.33, 19, 39, 40).
-While ANNs were previously used as value function approximators [29], the introduction of Deep Q-Networks (DQN) by [24, 25] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
-The experience replay (see 22) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation.
+Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.37, 20, 43, 44).
+While ANNs were previously used as value function approximators [32], the introduction of Deep Q-Networks (DQN) by Mnih et al.
+[25, 26] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
+Experience replay (see 23) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation.
 The target network further stabilizes the learning process by periodically copying the parameters of the learning network.
-This strategy is crucial because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence.
+This strategy is necessary because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence.
 By leveraging the target network, gradient steps are directed towards a periodically fixed target, ensuring more stability in the learning process.
-Additionally, the learning rate hyperparameter controls the magnitude of these gradient steps in optimizers such as the stochastic gradient descent algorithm, affecting the training convergence.
+Additionally, the learning rate [^1] The terms “learning rate” and “step-size” are used interchangeably in the literature and they technically refer to the same concept.
+hyperparameter controls the magnitude of these gradient steps in optimizers such as the Stochastic Gradient Descent (SGD) algorithm, affecting the training convergence.
 The learning rate is one of the most important hyperparameters, with previous work demonstrating that decreasing its value during policy finetuning can enhance performance by up to 25% in vanilla DQN [3].
-Determining the appropriate learning rate [^1] The terms “learning rate” and “step-size” are often used interchangeably in the literature and they technically refer to the same concept.
-is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 14, 7, 42).
+Determining the appropriate learning rate is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 15, 7, 46).
 However, finding a learning rate value that improves the model performance requires extensive and computationally expensive testing.
 In order to adapt its initial choice during training, optimizers such as Adam [18] and RMSProp [41] employ an internal scheme that dynamically adjusts the learning rate, considering, for instance, past gradient information.
-Nevertheless, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
+Therefore, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
 Standard learning rate schedulers typically decrease the learning rate based on training progress using, e.g., linear or exponential decay strategies [35, 46].
-In the context of RL, this approach can lead to premature or delayed learning rate adjustments, which may hinder the agent’s ability to learn.
 Unlike supervised learning, RL usually involves generating data by trading off between exploration (discovery of new states) and exploitation (refining of the agent’s knowledge).
@@
 Consequently, a more sophisticated decaying learning rate strategy that accounts for policy performance rather than training steps can significantly enhance learning in deep RL.
+Contributions.
 In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a method to select the learning rate on the fly for deep RL.
 Our approach acknowledges that different learning phases require different learning rates, and as such, instead of scheduling the learning rate decay using some blanket approach, we dynamically choose the learning rate using a Multi-Armed Bandit (MAB) algorithm, which accounts for the current policy’s performance.
-Our method has the advantage of being algorithm-agnostic and applicable to any optimizer, although the results show that it works best when coupled with Adam.
-We conduct experiments on our approach using baselines provided in the Dopamine framework (see 9).
-Our results focus on exploiting different settings for LRRL to illustrate its robustness under many possible configurations.
-Our main contributions are the following:
-  • We introduce LRRL, the first approach, to our knowledge, that leverages a multi-armed bandit algorithm to select the learning rate dynamically in deep RL.
-Our results demonstrate that LRRL achieves competitive performance with or superior to standard deep RL algorithms using fixed baselines or traditional learning rate schedulers.
-  • Our results show that LRRL significantly reduces the need for hyperparameter optimization by dynamically selecting from a set of possible learning rates using a multi-armed bandit approach.
-This method mitigates the need for exhaustive techniques like grid search, as it efficiently adapts the learning rate during training in a single run, rather than requiring multiple runs to test each learning rate individually.
-  • We assess the robustness of our method by employing the Adam and RMSProp optimizers with different sets of arms.
-We also compare the results using stochastic and adversarial multi-armed bandit algorithms in Appendix A.2.
-2 Related Work
-Multi-armed bandit for (hyper)parameter selection.
-Deep RL is known to be overly optimistic in the face of uncertainty [27], and many works have addressed this issue by proposing conservative policy updates [38, 13, 2].
-However, when the agent is able to interact with the environment, this optimism can encourage exploration, potentially leading to the discovery of higher returns.
-Building on this idea, [26] use an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process.
-In order to select the learning rate for stochastic gradient MCMC, [10] employ an algorithm based on Successive Halving [16, 15], a MAB strategy that promotes promising arms and prunes suboptimal ones over time.
-In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [21, 32].
-A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space.
-Close to our work, [23] propose Adam with Bandit Sampling (ADAMBS), which employs the Exponential-weight algorithm for Exploration and Exploitation [6] to enhance sample efficiency by incorporating importance sampling within the Adam optimizer.
-While ADAMBS prioritizes informative samples, our method focuses on RL tasks by dynamically adjusting the learning rate, accelerating learning when the policy is far from optimal, and slowing it down as it converges.
-Learning rate adapters/schedulers.
-Optimizers such as RMSProp [37] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [17] also incorporates momentum to smooth gradient steps.
-However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see  12).
-Increment-Delta-Bar-Delta (IDBD), introduced by  [34], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [43].
-Learning rate schedulers with time decay [31, 42] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning.
-A meta-gradient reinforcement learning is proposed in [41], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function.
-Our work differs from these methods by employing a multi-armed bandit approach to dynamically select the learning rate over the training process, specifically targeting RL settings.
+Our method has the advantage of being algorithm-agnostic and applicable to any optimizer.
+By exploiting different configurations, we aim to illustrate the use cases of LRRL and assess its robustness.
+We provide experiments with the Adam and RMSProp optimizers, with different sets of arms, and two deep RL algorithms (DQN and IQN).
+Tests are carried out across a variety of Atari games, using baselines provided in the Dopamine framework (see 9).
+We also show the effectiveness of LRRL using a standard SGD optimizer in a non-RL optimization context.
+Related work is detailed in Appendix A.
+Our main contributions and results can be summarized as follows:
+  • We introduce LRRL, a novel approach that leverages a MAB algorithm to dynamically select the learning rate during training in deep RL.
+  • We assess LRRL robustness across a set of Atari games, using different configurations, optimizers and deep RL algorithms.
+Our empirical results demonstrate that LRRL achieves performance competitive with or superior to standard learning using fixed learning rates or decay schedulers.
+We further highlight its advantages by considering an SGD optimizer on stationary non-convex landscapes.
+  • More generally, we show that LRRL matches the scheduler’s behavior by dynamically choosing lower learning rate values as the training progresses.
 2 Preliminaries
-This section introduces the Reinforcement Learning and Multi-Arm Bandits frameworks, defining supporting notation.
+This section introduces the RL and MAB frameworks, defining supporting notation.
 2.1 Deep Reinforcement Learning
-An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.35 for details).
-In RL, starting from an initial state $s_{0}$, a learner called agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$.
-In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(s_{t},a_{t},\,\cdot\,)$.
-The agent’s objective is to learn a policy $\pi:\mathcal{S\times A}\rightarrow[0,1]$ which maps a distribution of actions given the current state, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr{]}$.
-To learn how to perform a task, value function-based algorithms coupled with ANNs [24, 25] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$.
-By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function $\mathcal{J}(\theta)$ defined by:
+An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.39 for details).
+In RL, starting from an initial state $s_{0}$, a learner called the agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$.
+In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(\cdot|s_{t},a_{t})$.
+The agent’s objective is to learn a policy $\pi(\cdot|s)$ that maps a state $s$ to a probability distribution over actions in $\mathcal{A}$, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr{]}$.
+To learn how to perform a task, value function-based algorithms coupled with ANNs [25, 26] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(\cdot|s)=1_{a^{*}(s)}(\cdot)$ where $1_{a^{*}(s)}$ denotes the uniform distribution over ${a^{*}(s)}$ with $a^{*}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$.
+By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function defined by:
 $\mathcal{J}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\Bigl{[}r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})-Q_{\theta}(s,a)\Bigr{]}^{2}\,,$ (1)
@@
 The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$.
-The minimization, and hence, the update of the parameters $\theta$, is done according to the optimizer’s routine.
-A simple possibility is to use stochastic gradient descent using mini-batch approximations of the loss gradient:
-$\theta_{n+1}\leftarrow\theta_{n}-\eta\nabla_{\theta}\mathcal{J}(\theta_{n})\,,\quad\text{where}\quad\nabla_{\theta}\mathcal{J}(\theta_{n})\approx\frac{1}{|\mathcal{B}|}\sum_{(s,a,r,s^{\prime})\in\mathcal{B}}2\bigl{(}Q_{\theta}(s,a)-y\bigr{)}\nabla_{\theta}Q_{\theta}(s,a)$ (2)
+The minimization, and hence the update of the parameters $\theta$, is done according to the optimizer’s routine.
+A simple possibility is to use SGD using mini-batch approximations of the loss gradient:
+$\begin{gathered}\theta_{n+1}\leftarrow\theta_{n}-\eta\nabla_{\theta}\mathcal{J}(\theta_{n})\,,\text{where}\\ \nabla_{\theta}\mathcal{J}(\theta_{n})\approx\frac{1}{|\mathcal{B}|}\sum_{(s,a,r,s^{\prime})\in\mathcal{B}}2\bigl{(}Q_{\theta}(s,a)-y\bigr{)}\nabla_{\theta}Q_{\theta}(s,a),\end{gathered}$ (2)
 with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called the learning rate.
 Unlike in supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time.
-This continuous evolution introduces instability during the learning process, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$.
+This continuous evolution introduces instability during learning, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$.
 However, stability also depends on how the parameters $\theta$ change during each update.
-This work aims to control these changes by dynamically selecting the learning rate $\eta$ over the training steps.
-3.2 Multi-Armed Bandit
-Multi-Armed Bandits (MAB) provide an elegant framework for making sequential decisions under uncertainty (see for instance 20).
+This work aims to adapt to these changes by dynamically selecting the learning rate $\eta$ over the training steps.
+2.2 Multi-Armed Bandits
+MAB provide an elegant framework for making sequential decisions under uncertainty (see for instance 21).
 MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$.
 Like RL, MAB algorithms must balance the trade-off between exploring arms that have been tried less frequently and exploiting arms that have yielded higher rewards up to time $n$.
-To account for the non-stationarity of the RL rewards, we will consider in this work the MAB setting of adversarial bandits [6].
+To account for the non-stationarity of RL rewards, we will consider in this work the MAB setting of adversarial bandits [6].
 In this setting, at each round $n$, the agent selects an arm $k_{n}$ according to some distribution $p_{n}$ while the environment (the adversary) arbitrarily (e.g., without stationary constraints) determines the rewards $f_{n}(k)$ for all arms $k\in\mathcal{K}$.
 MAB algorithms are designed to minimize the pseudo-regret $G_{N}$ after $N$ rounds defined by:
-$G_{N}=\max_{k\in\{1,\dots,K\}}\,\mathbb{E}\Biggl{[}\sum_{n=1}^{N}f_{n}(k)-\sum_{n=1}^{N}f_{n}(k_{n})\Biggr{]}\,,$
-where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}f_{n}(k)}$ represents the accumulated reward of the single best arm in hindsight, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm.
+$G_{N}=\mathbb{E}\Biggl{[}\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)-\sum_{n=1}^{N}f_{n}(k_{n})\Biggr{]}\,,$
+where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)}$ represents the accumulated reward of the (pontentially changing) single best arm, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm.
 A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$: this guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
 3 Dynamic Learning Rate for Deep RL
-In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic Learning Rate for deep Reinforcement Learning (LRRL).
+In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic LRRL.
 LRRL is a meta-learning approach designed to dynamically select the learning rate in response to the agent’s performance.
-LRRL couples with stochastic gradient descent optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm.
+LRRL couples with SGD optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm.
 As the agent interacts with the environment, the average of observed rewards is used as bandit feedback to guide the selection of the most appropriate learning rate throughout the training process.
@@
 As we expect that the agent’s performance —and hence, the cumulative rewards— will improve over time, the MAB algorithm should receive non-stationary feedback.
-To take this non-stationary nature of the learning into account, we employ the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction).
+To take this non-stationary nature of the learning into account, we employ a modified version of the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction).
 At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively.
-Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to improvements in policy performance.
+Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to changes of the best action and to improvements in policy performance.
 Specifically, after picking arm $k_{n}$ at round $n$, the RL agent interacts $C$ times with the environment and the MAB algorithm receives a feedback $f_{n}$ corresponding to the average reward of those $C$ interactions.
-Based on [26], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
+Based on Moskovitz et al.
+[27], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
 $f_{n}^{\prime}=f_{n}-\frac{1}{j}\sum_{i=0}^{j-1}f_{n-i}\,.$
@@
 $\forall k\in\{1,\ldots,K\},\quad p_{n+1}(k)=\frac{e^{w_{n+1}(k)}}{\sum_{k^{\prime}=1}^{K}e^{w_{n+1}(k^{\prime})}}\,.$ (4)
-This update rule ensures that as the policy $\pi_{\theta}$ improves, the MAB algorithm continues to favor learning rates that are most beneficial under the current policy performance, thereby effectively handling the non-stationarity inherent in the learning process.
+This update rule ensures that as the policy $\pi_{\theta}$ improves the cumulative reward, the MAB algorithm continues to favor learning rates that are most beneficial under the current policy performance, thereby effectively handling the non-stationarity inherent in the learning process.
 4 Experiments
-In the following sections, we investigate whether combining LRRL with Adam or RMSProp —two widely used optimizers— can improve cumulative returns in deep RL algorithms.
-To assess this, we compare LRRL against learning methods with and without schedulers using the baseline implementation of DQN provided in Dopamine [9].
-We test LRRL under different configurations and Atari games, reporting the average and standard deviation of returns over 5 runs.
-Details on the evaluation metrics and hyperparameters used in these experiments are summarized in Appendix B.
+In the following sections, we examine whether combining LRRL with Adam or RMSProp —two widely used optimization algorithms— can enhance cumulative returns in deep RL algorithms.
+We compare LRRL against learning methods both with and without schedulers, using baseline algorithms provided in the Dopamine [9] framework.
+The experiments involve testing LRRL across various configurations and a selection of Atari games, reporting the mean and one-half standard deviation of returns over 5 independent runs.
+As testing the full Atari benchmark is extremely costly, we choose a representative subset of games guided by their baseline learning curves, which offer insights into reward sparsity, stochasticity, and exploration demands.
+Further details on the evaluation metrics and hyperparameters used in these experiments are provided in Appendices C and D.
+In addition, although our primary target is RL tasks, we also test LRRL using SGD to minimize stationary objective functions to provide a simpler evaluation, as performance in RL is highly multi-factorial.
 4.1 Comparing LRRL with Standard Learning
-In our first experiment, we consider a set of $5$ learning rates, and compare the performance of 5 configurations of LRRL, each of them using a subset of those learning rates, against the DQN algorithm reaching best performance, in terms of maximum average return, among the $5$ possible learning rates choices (see Figure 5 in Appendix A.1).
-More precisely, the set of learning rates is
+In our first experiment, we consider a set of $5$ learning rates.
+We set different configurations of LRRL using these values and compare them against the DQN agent reaching the best performance, in terms of maximum average return, among the possible single rates.
+More specifically, the learning rates considered are
 $\mathcal{K}(5)=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}\,.$
-The whole set $\mathcal{K}(5)$ is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) and 3 taking the lowest/middle/highest ($\mathcal{K}_{\text{sparse}}(3)$) learning rate values.
+The whole set $\mathcal{K}(5)$ of rates is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) values, and the last one ($\mathcal{K}_{\text{sparse}}(3)$) taking the lowest/middle/highest values.
 All experiments use the Adam optimizer, and we report the return based on the same number of environment iterations.
-Results are gathered in Table 1 and illustrated in Figure 1 along with four Atari games.
-They show that LRRL outperforms standard learning in two out of four tasks while remaining competitive in the others.
-Notably, LRRL not only matches or exceeds performance but also reduces the need for extensive parameter tuning by incorporating multiple learning rates in a single run.
-However, as the results indicate, the $3$-arm bandit variants exhibit different behaviors during the learning process, suggesting that the choice of the number of arms and their values still requires task-specific tuning to achieve good performance.
-Figure 1: A comparison between 5 configurations of LRRL with various learning rate subsets and the DQN algorithm reaching the best performance among possible learning rates.
-Game DQN LRRL $\mathcal{K}_{\text{lowest}}(3)$ LRRL $\mathcal{K}_{\text{middle}}(3)$ LRRL $\mathcal{K}_{\text{highest}}(3)$ LRRL $\mathcal{K}_{\text{sparse}}(3)$ LRRL $\mathcal{K}(5)$
-Asteroids 1 085 $\pm$ 63 1 065 $\pm$ 70 1 079 $\pm$ 48 1 061 $\pm$ 59 1007 $\pm$ 127 1 085 $\pm$ 88
-Breakout 217 $\pm$ 14 220 $\pm$ 24 201 $\pm$ 6 177 $\pm$ 8 270 $\pm$ 13 253 $\pm$ 20
-Pong 19 $\pm$ 0.4 18 $\pm$ 0.3 19 $\pm$ 0.9 19 $\pm$ 0.5 19 $\pm$ 0.5 19 $\pm$ 0.8
-Seaquest 5 881 $\pm$ 1 533 5 905 $\pm$ 2 557 6 135 $\pm$ 2 229 6 231 $\pm$ 1 802 8 920 $\pm$ 2 759 6 799 $\pm$ 2 060
-Table 1: Max average return (best in bold if significantly better than others) and its standard deviation for 4 Atari games.
-To illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves, Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$.
+The results in Figure 1 demonstrate that LRRL using the full set $\mathcal{K}(5)$ substantially outperforms the best-performing single learning rate in 2 out of 6 games (Seaquest and Breakout) while remaining competitive in the others.
+As shown in Figure 7 in Appendix B.1, the baseline DQN performance is highly sensitive to the learning rate choice, requiring extensive hyperparameter tuning to identify the optimal rate.
+In contrast, LRRL dynamically selects learning rates during training, avoiding the need for such tuning.
+Additionally, the $3$-arm bandit variants highlight that while LRRL can explore multiple learning rates within a single run, the specific set of rates available still impacts performance.
+Overall, these results illustrate the advantage of LRRL, which achieves similar or better performance than the best-tuned DQN baseline while eliminating the computational cost of manual learning rate tuning.
+Figure 1: Comparison between LRRL and standard DQN learning: 5 variants of LRRL with different learning rate sets (provided in Appendix D) are tested again the DQN algorithm reaching best performance among the possible learning rates.
+The mean and one-half standard deviations over 5 runs are respectively represented by the curves and shaded areas.
+Next, we illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves.
+Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$.
 In most of the tested environments, LRRL behaves similarly to time-decay schedulers by selecting higher learning rates during the early stages of training, gradually shifting toward arms with lower rates as training progresses.
-The exception is in the Pong environment, where the model converges after only a few iterations, resulting in a more uniform probability distribution across the set of arms.
 Figure 2: Systematic sampling of normalized learning rates and returns over the training steps using LRRL $\mathcal{K}(5)$ with Adam optimizer, through a single run.
 For each episode, we show the selected learning rate using different colors.
+Lower rates are increasingly selected over time.
 4.2 Combining and Comparing Schedulers with LRRL
-Next, we consider using LRRL combined with learning rate schedulers.
+In the following experiments, LRRL is combined with learning rate schedulers.
 Specifically, we employ schedulers with exponential decay rate of the form $\eta(n)=\eta_{0}\times e^{-d\,n}$, where $\eta_{0}$ is a fixed initial value (common to each scheduler and equal to $6.25\times 10^{-5}$ in our experiment), $d$ is the exponential decay rate and $n$ is the number of policy updates (i.e., of MAB rounds).
 We define a set of $3$ schedulers $\mathcal{K}_{s}$, where each arm represents a scheduler using a different decay rate $d=\{1,2,3\}\times 10^{-7}$, and compare the results of LRRL with each scheduler individually, using the Adam optimizer.
-Figure 3 and Table 2 show that LRRL combined with schedulers can substantially increase final performance compared to using exponential decay schedulers for some environments while remaining competitive for others.
-The dashed black line represents the max average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by [4], who linearly decay the learning rate to 00.
-Figure 3: A comparison between LRRL with arms as schedulers and schedulers individually.
-Game DQN $d=1\times 10^{-7}$ $d=2\times 10^{-7}$ $d=3\times 10^{-7}$ LRRL ($\mathcal{K}_{s}$)
-Asteroids 1 028 $\pm$ 53 1 025 $\pm$ 52 1 013 $\pm$ 56 1 063 $\pm$ 82 1 015 $\pm$ 33
-Breakout 144 $\pm$ 12 149 $\pm$ 11 151 $\pm$ 7 144 $\pm$ 12 233 $\pm$ 19
-Seaquest 5 881 $\pm$ 1 533 5 284 $\pm$ 1 134 5 793$\pm$ 1 225 6 612 $\pm$ 835 13 864 $\pm$ 3 581
-Video Pinball 410 186 $\pm$ 193 328 388 768 $\pm$ 195 150 450 015 $\pm$ 178 876 362 645 $\pm$ 172 169 388 308 $\pm$ 103 862
-Table 2: Max average return (best in bold) and its standard deviation for 4 Atari games.
-5.3 RMSProp Optimizer and more Environments
-Another widely used optimizer for training deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism.
+Figure 3 shows that LRRL combined with schedulers could once again substantially increase the final performance of 2 out of 6 tasks compared to using a fixed scheduler.
+In the remaining tasks, LRRL remains at least competitive, while for most of them, it shows a better jumpstart performance.
+The dashed black line represents the maximal average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by Andrychowicz et al.
+[4], who linearly decay the learning rate to 00.
+In all the experiments conducted so far, we have set the update window $\kappa$ parameter to 1, which allows the observation of a full episode before updating the learning rate.
+In Appendix B.3, we address how this value could influence the performance of LRRL by allowing the observation of more interactions.
+Figure 3: LRRL with arms that select learning rate schedulers: LRRL vs 3 individual schedulers ($d=1,2,3$).
+The dashed black line represents the max average return achieved by Adam with a constant learning rate set to $6.25\times 10^{-5}$.
+4.3 Comparing RMSProp and Adam Optimizers
+Another widely used optimizer for deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism.
 Adam builds upon RMSProp by retaining exponential moving averages to give more weight to recent gradients while incorporating momentum.
@@
 In the following experiment, we compare the performance of RMSProp with Nesterov’s momentum (RMSProp-M), and Adam when coupled with either LRRL or the best-performing single learning rate when using DQN.
-As shown in Figure 4 and Table 3, LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN.
+Since the results were consistent in previous sections, we extend the experiments to other 3 Atari games.
+Figure 4 shows that LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN.
 Moreover, LRRL (RMSProp-M) underperforms compared to DQN without LRRL in two out of three tasks due to its slow convergence despite better jumpstart performance.
-Future work should investigate whether this slow convergence is linked to factors such as the environment’s stochasticity or the optimizer’s features such as the absence of bias correction in the first and second moment estimates.
+Future work should investigate whether this slow convergence is due to some optimizer features such as the absence of bias correction in the first and second moment estimates.
 Figure 4: A comparison between Adam and RMSProp with momentum, using either DQN or LRRL.
-Game DQN (RMSProp-M) LRRL (RMSProp-M) DQN (Adam) LRRL (Adam)
-Asterix 11 464 $\pm$ 2 848 6 499 $\pm$ 927 12 561 $\pm$ 1 245 15 017 $\pm$ 3 892
-Ms. Pacman 3 301 $\pm$ 310 2 696 $\pm$ 248 3 232 $\pm$ 114 3 310 $\pm$ 231
-Space Invaders 1 490 $\pm$ 132 2 712 $\pm$ 73 2 874 $\pm$ 319 3 641 $\pm$ 1 030
-Table 3: Max average return (best in bold) and its standard deviation for 3 Atari games.
+4.4 LRRL with a Distributional RL Algorithm
+Lastly, we illustrate the versatility of LRRL by combining it with the distributional RL algorithm Implicit Quantile Networks (IQN) [11], which exhibits the best performance in many Atari games among the algorithms available in Dopamine.
+Adam is used as the optimizer.
+Figure 5 shows that LRRL outperforms standard learning with a single learning rate across all the tasks evaluated.
+More generally, LRRL can be seamlessly integrated with any RL algorithm that employs function approximators based on SGD, improving its performance and mitigating hyperparameter tuning.
+Figure 5: Comparison between LRRL and standard IQN learning: IQN combined with LRRL using 3 different learning rates is tested against the 3 variants of standard IQN using each possible learning rate.
+4.5 SGD and Stationary Non-convex Loss Landscapes
+Although RL tasks are a challenging showcase to illustrate the effectiveness of our approach in the non-stationary case, there are many confounding factors to analyze their impact on the results, such as the optimizer adaptive routine, the environment stochasticity, and the reward sparseness.
+To disentangle these factors from our evaluation, in the following experiments we employ the SGD optimizer, which has a single scalar learning rate without adaptive scheme or momentum to find the global minima of $6$ non-convex loss landscapes.
+Since these functions represent a stationary case, we employ the stochastic bandit Minimax Optimal Strategy in the Stochastic case (MOSS) algorithm (see details in Appendix B.2), using as feedback the loss plus a small constant $\xi$, such that $f_{n}=\frac{1}{|\textrm{loss}(n)|+\xi}$, observed at each gradient step $n$ to update the arm’s probabilities.
+Figure 6 illustrates the performance of LRRL against SGD with a single learning rate, starting from the same initial point and for the same amount of steps.
+The final steps are indicated by a star icon with a different color.
+The results show that LRRL, with its “all-in-one” strategy, efficiently navigates intricate landscapes.
+It is effective in combining higher learning rates to accelerate convergence and lower ones for more fine-grained optimization steps when needed, as evidenced in the loss landscapes depicted in Figures 6(a)6(b), and 6(c).
+In addition, Figures 6(d)6(e), and 6(f) illustrate that LRRL can be also effective in converging towards the best-performing learning rate within the arms, showcased by its steady convergence towards the global minima.
+Figure 6: Comparison between LRRL and fixed learning rates in a stationary setting with SGD.
+LRRL with 3 learning rates is tested against the 3 constant rates for 6 non-convex losses (shown in Figure 10 in Appendix B.4).
+LRRL is effective in either converging to the best performing rate or combining the set to achieve better results.
 5 Conclusion
-In this work, we introduced dynamic Learning Rate for Deep Reinforcement Learning (LRRL), a meta-learning approach for selecting the optimizer’s learning rate on the fly.
-We demonstrated empirically that combining LRRL with the Adam optimizer could significantly enhance the performance of the value-based algorithm DQN, outperforming baselines and learning rate schedulers in some tasks while remaining competitive in others.
-Furthermore, by employing a multi-armed bandit algorithm, LRRL reduces the need for extensive hyperparameter tuning, as it explores a set of learning rates in a single run with minimal extra computational overhead.
-While this work focused on dynamically selecting the best-performing learning rate, future investigations could extend LRRL ideas to other critical hyperparameters, such as mini-batch size, which also plays a key role in the model’s convergence.
-Moreover, although LRRL selects learning rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [30, 18, 1].
+In this work, we introduced LRRL, a meta-learning approach for selecting optimizers’ learning rates on the fly.
+In any optimization task, the best-performing learning rate or rate scheduler is typically unknown, requiring multiple runs to test each candidate in a wasteful computational effort.
+LRRL mitigates this cost by employing a multi-armed bandit algorithm to explore a set of learning rates within a single run, significantly reducing the need for extensive hyperparameter tuning while achieving better or competitive results with minimal computational overhead.
+Using RL and non-convex optimization tasks, our empirical results show that combining LRRL with the Adam or SGD optimizers could enhance the performance, outperforming baselines and learning rate schedulers.
+Future investigations could extend LRRL ideas to other hyperparameters, such as mini-batch size, which also plays a key role in convergence.
+Moreover, although LRRL selects rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [33, 19, 1].
+Limitations.
+The primary limitation of LRRL lies in the need to define a set of arms for the MAB algorithm, as the choice of this set can directly influence the performance of the underlying learning algorithm.
+However, in computationally intensive settings such as RL tasks, LRRL offers a significant advantage: it enables the simultaneous evaluation of multiple learning rates within a single training run, eliminating the need to test each rate individually.
+By employing an “all-in-one” strategy, LRRL not only achieves competitive results while reducing the burden of hyperparameter optimization but also can outperform standard learning by combining multiple learning rates.
+Broader Impact Statement
+This work aims to contribute to the advancement of the field of Machine Learning.
+While we did not identify any direct negative social impacts associated with this research, we encourage future studies building on it to remain vigilant and critically evaluate potential ethical and societal implications.
 Acknowledgment
@@
 LFS is supported by a Discovery Early Career Researcher Award from the Australian Research Council (DE240101190)
-Appendix A Supplementary Experiments
-A.1 Baseline Evaluation with Varying Learning Rates
+Appendix A Related Work
+This section presents works closely related to ours.
+We first introduce approaches that use MAB for hyperparameter selection and later other methods that aim to adapt the learning rate over the training process.
+Multi-armed bandit for (hyper)parameter selection.
+Deep RL is known to be overly optimistic in the face of uncertainty [28], and many works have addressed this issue by proposing conservative policy updates [42, 14, 2].
+However, when the agent can interact with the environment, this optimism might encourage exploration, potentially leading to the discovery of higher returns.
+Building on this idea, Moskovitz et al.
+[27] proposes Tactical Optimism and Pessimism (TOP), an approach that uses an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process.
+Although our work shares similarities with TOP, such as employing an adversarial bandit algorithm and using performance improvement as feedback, there are key differences.
+First, TOP is closely tied to actor-critic methods, making it unclear how it could be extended to other RL paradigms.
+Moreover, it also introduces an exploration strategy which depends on epistemic uncertainty (captured through ensembles) and aleatoric uncertainty (estimated using a distributional RL algorithm).
+Second, TOP trades-off between optimism and pessimism by adding new variables (arms), which are not inherently tied to the learning process and also require tuning.
+In contrast, LRRL offers an algorithm-agnostic solution by directly optimizing the learning rate, a core hyperparameter in all SGD-based methods, making it simpler and more broadly applicable.
+Also close to our work, Liu et al.
+[24] propose Adam with Bandit Sampling (ADAMBS), which employs the Exp3 algorithm to enhance sample efficiency by incorporating importance sampling within the Adam optimizer.
+ADAMBS prioritizes informative samples by keeping a probability distribution over all samples, although the feedback is provided by the gradient computed over a mini-batch, unlike previous works in RL that are capable of assigning per-sample importance by using their computed TD-error (see 34).
+In order to select the learning rate for stochastic gradient MCMC, Coullon et al.
+[10] employ an algorithm based on Successive Halving [17, 16], a MAB strategy that promotes promising arms and prunes suboptimal ones over time.
+In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [22, 36].
+A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate on the fly from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space.
+Close to our work, Parker-Holder et al.
+[29] propose population-based training which employs a Gaussian Process (see details in 31) with a time-varying kernel to represent the upper confidence bound in a bandit setting.
+The idea is to have many instances of agents learning a task in parallel with different hyperparameters, replacing those that are underperforming within a single run.
+Learning rate adapters/schedulers.
+Optimizers such as RMSProp [41] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [18] also incorporates momentum to smooth gradient steps.
+However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see 13).
+Increment-Delta-Bar-Delta (IDBD), introduced by Sutton [38], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [47].
+Learning rate schedulers with time decay [35, 46] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning.
+A meta-gradient RL method is proposed in Xu et al.
+[45], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function.
+Our work differs from these methods by employing a MAB approach to dynamically select the learning rate over the training process, specifically targeting RL settings.
+Appendix B Supplementary Experiments
+B.1 Baseline Evaluation with Varying Learning Rates
 To establish a baseline to compare learning without our approach LRRL, we run individual arm values as baseline learning rate using the Adam optimizer.
-The results presented in Figure 5 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to the worst jumpstart performance and slow convergence.
-Figure 5: DQN performance using Adam optimizer with varying learning rates across 4 Atari games.
-A.2 A Comparison between Multi-Armed Bandits Algorithms
+The results presented in Figure 7 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to worse jumpstart performance and slow convergence.
+Figure 7: DQN performance using the Adam optimizer with varying learning rates across 6 Atari games.
+Shaded region represents one-half standard deviation.
+The results show that the best performing learning rate varies across games.
+B.2 A Comparison between Multi-Armed Bandits Algorithms
 Adversarial MAB algorithms are designed for environments where the reward distribution changes over time.
@@
 To validate our choice and address our method’s robustness, we compare Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [5].
-MOSS trade-off exploration-exploitation by pulling the arm $k$ with highest upper confidence bound given by:
+MOSS trades off exploration and exploitation by pulling the arm $k$ with highest upper confidence bound given by:
 $B_{k}(n)=\hat{\mu}_{k}(n)+\rho\,\sqrt{\frac{\max\Bigl{(}\log{\frac{n}{Kn_{k}(n)}},0\Bigr{)}}{n_{k}(n)}}$
-where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to timestep $t$.
-In Figure 6, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balance exploration and exploitation in MOSS.
+where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to round $n$.
+In Figure 8, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balances exploration and exploitation in MOSS.
 Additionally, the bandit feedback used in MOSS is the average cumulative reward $\frac{R}{C}$.
 The results indicate that while Exp3 performs better overall, MOSS can still achieve competitive results depending on the amount of exploration, demonstrating the robustness of our approach regarding the MAB algorithm employed.
-Figure 6: A comparison between adversarial and stochastic MAB algorithms.
-Appendix B Experiments Details
-B.1 Evaluation Terminology
-In this section, we describe the evaluation metrics that can be used to evaluate agent’s performance as it interacts with an environment.
-Based on Dopamine, we use the evaluation step-size “iterations”, which is defined as a predetermined number of episodes.
-Figure 7 illustrates the evaluation metrics used in this work, as defined in [36]:
+Figure 8: A comparison between adversarial (Exp3) and stochastic (MOSS) MAB algorithms. Shaded region represents one-half standard deviation. The results show that Exp3 performs better overall.
+B.3 Dealing with Noise and Lack of Feedback
+LRRL relies entirely on performance improvements to update the arm probabilities.
+However, some RL environments can be highly stochastic, with sparse or delayed rewards, which increases the difficulty of learning tasks and, consequently, the ability of LRRL to make optimal choices.
+To mitigate this limitation, we introduce the variable $\kappa$ in Algorithm 1, which determines the number of interactions to observe before making a decision.
+When setting $\kappa$, two factors should be considered:
+  • The network’s update rate: Selecting new learning rates more frequently than updating the model leads to inaccurate feedback for the bandit mechanism since no model update might occur after pulling some arm.
+  • The number of environment steps: In some tasks, positive or negative outcomes of a sequence of actions are only perceived at the end of an episode, and capturing sufficient information about the agent performance is essential to updating the arms’ probabilities.
+In the experiments presented in this work, we set $\kappa=1$ episode, allowing the observation of a complete episode before updating the arms’ probabilities.
+To explore the impact of this parameter further, we conduct an ablation study using different $\kappa$ values to evaluate potential performance improvements when combined with the schedulers shown in Figure 3.
+As suggested by Figure 9, the choice of $\kappa$ is task-dependent: while some tasks benefit from more reactive learning rate updates, others show marginal improvements with increased observation counts.
+Ultimately, setting $\kappa=1$ seems relatively robust for Atari games, although higher values could slightly increase the performance in some tasks.
+Figure 9: Results of the ablation study evaluating the impact of varying $\kappa$ with LRRL using schedulers.
+The dashed black line represents the max average return achieved by the best performing scheduler.
+B.4 Supplementary Results using SGD
+In Section 4.5, we tested the SGD optimizer in $6$ non-convex loss landscapes show in Figure 10.
+These functions were selected since they represent common benchmarks in optimization, each one offering different challenges due to their distinct shapes.
+By using SGD on these landscapes, we minimize the influence of confounding factors, such as the high stochasticity inherent in some RL tasks and the complexity of more robust optimizers like Adam.
+Figure 11 presents the normalized loss and the arms selected at each optimization step.
+The results demonstrate LRRL’s dynamic adaptation mechanism, either converging to the most effective learning rate or combining them to achieve superior outcomes.
+Figure 10: The 6 stationary non-convex loss landscapes used to assess LRRL with SGD optimizer.
+Figure 11: Learning rates and logarithmic scale of losses over the gradient steps using LRRL with SGD optimizer.
+Appendix C Evaluation Metrics
+In this section, we describe the evaluation metrics that can be used to evaluate the agent’s performance as it interacts with an environment.
+Based on Dopamine, we use the evaluation step-size “iterations”, defined as a predetermined number of episodes.
+Figure 12 illustrates the evaluation metrics used in this work, as defined in Taylor and Stone [40]:
   • Max average return: The highest average return obtained by an algorithm throughout the learning process. It is calculated by averaging the outcomes across multiple individual runs.
@@
   • Jumpstart performance: The performance at the initial stages of training, starting from a policy with randomized parameters $\theta$.
-In Figure 7, Algorithm B exhibits better jumpstart performance but ultimately achieves lower final performance than Algorithm A.
+In Figure 12, Algorithm A exhibits better jumpstart performance but ultimately achieves lower final performance than Algorithm B.
 A lower jumpstart performance can result from factors such as a lower learning rate, although this work demonstrates that this does not necessarily lead to worse final performance.
-Figure 7: Performance curves of two RL algorithms (adapted from taylor2009transfer [36]).
-B.2 Further Experimental Details
+Figure 12: Performance curves of two RL algorithms (adapted from 40).
+Appendix D Implementation Details
 In the following, we list the set of arms, optimizer and the bandit step-size used in each experiment.
@@
     $\mathcal{K}_{\text{lowest}}(3)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
-    $\hskip 25.03856pt\mathcal{K}_{\text{middle}}(3)=\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
-    $\hskip 84.50471pt\mathcal{K}_{\text{highest}}(3)=\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
+    $\mathcal{K}_{\text{middle}}(3)$ =$=\hskip 65.44142pt\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
+    $\mathcal{K}_{\text{highest}}(3)$ =$=\hskip 125.19194pt\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
 […]
@@
   • Initial learning rate of schedulers: $\eta_{0}=6.25\times 10^{-5}$
-Section 5.3 – RMSProp Optimizer and More Environments.
-  • Optimizer: RMSProp for baselines
+Section 4.3– Comparing RMSProp and Adam Optimizers
+  • Optimizer: RMSProp for DQN baselines
   • Bandit step-size: $\alpha=0.2$
-  • LRRL set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
-Section A.2 – A Comparison between Multi-Armed Bandit Algorithms.
+  • LRR-DQN set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
+Section 4.4– LRRL with a Distributional RL Algorithm.
+  • Optimizer: Adam
+  • Bandit step-size: $\alpha=0.2$
+  • LRRL-IQN set of learning rates: $\mathcal{K}=\bigl{\{}2.5\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\bigr{\}}$
+Section B.2 – A Comparison between Multi-Armed Bandit Algorithms.
   • Optimizer: Adam
   • MOSS/Exp3 set of learning rates: $\mathcal{K}=\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
-B.3 Hyperparameters
-In this section, we outline the hyperparameters used in the experiments.
-The optimizers from the Optax library [11] are employed alongside the JAX [8] implementation of the DQN algorithm [24, 25], as provided by the Dopamine framework [9].
-Hyperparameter Setting
+Hyperparameters of the optimizers and deep RL algorithms.
+For reproductibility, the list of hyperparameters of the optimizers and deep RL algorithms used for our experiments is provided in Table 1.
+The optimizers from the Optax library [12] are employed alongside the JAX [8] implementation of the DQN [25, 26] and IQN [11] algorithms, as provided by the Dopamine framework [9].
+DQN and IQN hyperparameters Value
 Sticky actions True
@@
 Exploration decay rate 0.01
-Exploration decay period 250000 steps
-Environment steps per iteration 250000 steps
+Exploration decay period (steps) 250 000
+Environment steps per iteration (steps) 250 000
 Reward clipping [-1, 1]
 Network neurons per layer 32, 64, 64
-Hardware V100 GPU
+Hardware (GPU) V100
 Adam hyperparameters
@@
 $\beta_{2}$ decay 0.999
-Eps 1.5e-4
-RMSProp hyperparameters
+Eps (DQN) 1.5$\times$10-4
+Eps (IQN) 3.125$\times$10-4
+RMSProp hyperparameters (DQN)
 Decay 0.9
@@
 Centered False
-Eps 1.5e-4
-Table 4: Hyperparameters
+Eps 1.5$\times$10-4
+Table 1: Hyperparameters used in the experiments.

### v3

@@
 Abstract
-In deep Reinforcement Learning (RL) models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence.
-Additionally, due to the non-stationarity of the objective function, the best-performing learning rate can change over the training steps.
-To adapt the learning rate, a standard technique consists of using decay schedulers.
-However, these schedulers assume that the model is progressively approaching convergence, which may not always be true, leading to delayed or premature adjustments.
-In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that selects the learning rate based on the agent’s performance during training.
-LRRL is based on a multi-armed bandit algorithm, where each arm represents a different learning rate, and the bandit feedback is provided by the cumulative returns of the RL policy to update the arms’ probability distribution.
-Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms for some tasks.
+In deep Reinforcement Learning (RL), the learning rate critically influences both stability and performance, yet its optimal value shifts during training as the environment and policy evolve.
+Standard decay schedulers assume monotonic convergence and often misalign with these dynamics, leading to premature or delayed adjustments.
+We introduce LRRL, a meta-learning approach that dynamically selects the learning rate based on policy performance rather than training steps.
+LRRL adaptively favors rates that improve returns, remaining robust even when the candidate set includes values that individually cause divergence.
+Across Atari and MuJoCo benchmarks, LRRL achieves performance competitive with or superior to tuned baselines and standard schedulers.
+Our findings position LRRL as a practical solution for adapting to non-stationary objectives in deep RL.
 1 Introduction
-Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.37, 20, 43, 44).
-While ANNs were previously used as value function approximators [32], the introduction of Deep Q-Networks (DQN) by Mnih et al.
-[25, 26] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
-Experience replay (see 23) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation.
-The target network further stabilizes the learning process by periodically copying the parameters of the learning network.
-This strategy is necessary because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence.
-By leveraging the target network, gradient steps are directed towards a periodically fixed target, ensuring more stability in the learning process.
-Additionally, the learning rate [^1] The terms “learning rate” and “step-size” are used interchangeably in the literature and they technically refer to the same concept.
-hyperparameter controls the magnitude of these gradient steps in optimizers such as the Stochastic Gradient Descent (SGD) algorithm, affecting the training convergence.
-The learning rate is one of the most important hyperparameters, with previous work demonstrating that decreasing its value during policy finetuning can enhance performance by up to 25% in vanilla DQN [3].
-Determining the appropriate learning rate is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 15, 7, 46).
-However, finding a learning rate value that improves the model performance requires extensive and computationally expensive testing.
-In order to adapt its initial choice during training, optimizers such as Adam [18] and RMSProp [41] employ an internal scheme that dynamically adjusts the learning rate, considering, for instance, past gradient information.
-Therefore, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
-Standard learning rate schedulers typically decrease the learning rate based on training progress using, e.g., linear or exponential decay strategies [35, 46].
-Unlike supervised learning, RL usually involves generating data by trading off between exploration (discovery of new states) and exploitation (refining of the agent’s knowledge).
-As the policy improves, the data distribution encountered by the agent becomes more concentrated, but this evolution occurs at a different pace than the overall training progress.
-For instance, some environments require extensive exploration due to the sparseness of rewards, while others need more exploitation to refine the policy to the complexity of the task.
-Consequently, a more sophisticated decaying learning rate strategy that accounts for policy performance rather than training steps can significantly enhance learning in deep RL.
-Contributions.
-In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a method to select the learning rate on the fly for deep RL.
-Our approach acknowledges that different learning phases require different learning rates, and as such, instead of scheduling the learning rate decay using some blanket approach, we dynamically choose the learning rate using a Multi-Armed Bandit (MAB) algorithm, which accounts for the current policy’s performance.
-Our method has the advantage of being algorithm-agnostic and applicable to any optimizer.
-By exploiting different configurations, we aim to illustrate the use cases of LRRL and assess its robustness.
-We provide experiments with the Adam and RMSProp optimizers, with different sets of arms, and two deep RL algorithms (DQN and IQN).
-Tests are carried out across a variety of Atari games, using baselines provided in the Dopamine framework (see 9).
-We also show the effectiveness of LRRL using a standard SGD optimizer in a non-RL optimization context.
-Related work is detailed in Appendix A.
-Our main contributions and results can be summarized as follows:
-  • We introduce LRRL, a novel approach that leverages a MAB algorithm to dynamically select the learning rate during training in deep RL.
-  • We assess LRRL robustness across a set of Atari games, using different configurations, optimizers and deep RL algorithms.
-Our empirical results demonstrate that LRRL achieves performance competitive with or superior to standard learning using fixed learning rates or decay schedulers.
-We further highlight its advantages by considering an SGD optimizer on stationary non-convex landscapes.
-  • More generally, we show that LRRL matches the scheduler’s behavior by dynamically choosing lower learning rate values as the training progresses.
+Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.39, 21, 47, 48).
+These achievements largely depend on training neural networks with gradient-based optimizers, where the choice of learning rate is critical to stability and convergence.
+An overly large learning rate can lead to divergence, while a small one can result in prohibitively slow learning (see 17, 6, 50).
+Despite its importance, selecting an appropriate learning rate in RL remains challenging due to two intertwined factors: the non-stationarity of the objective and the high cost of hyperparameter tuning.
+Unlike supervised learning, where the loss function is typically stationary, RL agents learn from data generated by an evolving policy.
+As training progresses, both the distribution of visited states and the policy-induced rewards change, making the optimization landscape highly dynamic.
+Consequently, a learning rate that works well early in training may later become suboptimal.
+Prior work shows that even simple adjustments, such as reducing the learning rate in later stages, can improve performance in deep RL [2].
+However, designing an effective schedule is difficult because training steps do not necessarily reflect monotonic convergence toward a better policy.
+Decay schedules, such as linear or exponential decay [37, 50], reduce the learning rate over time based on predefined rules, assuming that the optimization problem becomes easier as training proceeds.
+This assumption often fails in RL, where progress can plateau or regress due to exploration challenges, sparse rewards, or environment stochasticity.
+Similarly, adaptive optimizers such as RMSProp [45] and Adam [20] adjust effective learning rates (see 25) based on past gradient statistics, which stabilizes updates but does not account for shifts in policy quality.
+Because these methods rely on step counts or gradient norms rather than environment feedback, they often make premature or delayed adjustments, limiting their ability to cope with RL’s non-stationarity.
+This motivates the question: Can we adapt the learning rate dynamically based on the agent’s performance, rather than relying on training progress or gradient-based heuristics?
+A method satisfying this criterion should (i) leverage feedback that reflects policy quality rather than indirect signals like gradient norms, (ii) adapt to the different dynamics of early exploration and late-stage refinement, and (iii) remain general enough to integrate with a wide range of algorithms and optimizers.
+We address this challenge by proposing Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that casts learning rate adaptation as a dual objective: while the RL agent seeks to maximize cumulative returns, the bandit algorithm minimizes regret in selecting the most effective learning rates over time.
+LRRL formulates this as an adversarial multi-armed bandit (MAB) problem, where each arm corresponds to a candidate learning rate, and feedback is based on policy performance.
+By leveraging adversarial bandit algorithms with time-decay weighting, LRRL balances exploration of alternatives with exploitation of promising rates.
+Unlike fixed schedules or gradient-based meta-learning heuristics, LRRL is algorithm-agnostic, seamlessly integrates with optimizers such as Adam, RMSProp, and SGD, and enables dynamic adaptation across different phases of training.
+We evaluate LRRL across a diverse set of scenarios, including discrete and continuous control tasks (Atari and MuJoCo), multiple deep RL algorithms (DQN, IQN, and PPO), and various optimizers.
+Our results show that LRRL achieves competitive or superior performance compared to fixed learning rates and standard decay strategies, while also demonstrating robustness to variations in configuration.
+Furthermore, to isolate RL-specific confounders, such as exploration and reward sparsity, we assess LRRL on stationary non-convex optimization problems using SGD, illustrating its ability to adaptively combine aggressive and conservative updates in a single optimization trajectory.
+Our main contributions are:
+  • We introduce LRRL, a novel approach that leverages a bandit-based controller to select among a small set of learning rates on the fly using policy performance as feedback.
+This directly targets RL’s non-stationarity and avoids assuming monotonic convergence.
+  • We demonstrate robustness to poorly performing candidates: LRRL remains stable and competitive even when the candidate set contains rates that individually fail, by adapting toward effective rates during training.
+  • We evaluate LRRL using a set of learning algorithms on Atari and MuJoCo tasks, showing performance competitive with or superior to tuned fixed-rate training and decay schedulers, alongside ablations of key design parameters.
+  • Beyond RL, we illustrate how combining multiple rates within a single run can be beneficial in stationary non-convex optimization with SGD, underscoring LRRL’s generality.
 2 Preliminaries
-This section introduces the RL and MAB frameworks, defining supporting notation.
+This section introduces the RL and MAB frameworks and defines the supporting notation.
 2.1 Deep Reinforcement Learning
-An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.39 for details).
-In RL, starting from an initial state $s_{0}$, a learner called the agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$.
+An RL task is defined by a Markov Decision Process (MDP), that is, by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see e.g.42 for details).
+In RL, starting from an initial state $s_{0}$, a learner called agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$.
 In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(\cdot|s_{t},a_{t})$.
@@
 To learn how to perform a task, value function-based algorithms coupled with ANNs [28, 29] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(\cdot|s)=1_{a^{*}(s)}(\cdot)$ where $1_{a^{*}(s)}$ denotes the uniform distribution over $a^{*}(s)$} with $a^{*}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$.
-By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function defined by:
+By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ in replay memory $\mathcal{D}$, the objective is to minimize the loss function defined by:
 $\mathcal{J}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\Bigl[r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})-Q_{\theta}(s,a)\Bigr]^{2}\,$, (1)
-where $\theta^{-}$ are the target network parameters used to compute the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$.
+where $\theta^{-}$ are the parameters of the target network used to calculate the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$.
 The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$.
@@
 $\begin{gathered}\theta_{n+1}\leftarrow\theta_{n}-\eta\nabla_{\theta}\mathcal{J}(\theta_{n})\,,\text{where}\\ \nabla_{\theta}\mathcal{J}(\theta_{n})\approx\frac{1}{|\mathcal{B}|}\sum_{(s,a,r,s^{\prime})\in\mathcal{B}}2\bigl(Q_{\theta}(s,a)-y\bigr)\nabla_{\theta}Q_{\theta}(s,a),\end{gathered}$ (2)
-with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called the learning rate.
-Unlike in supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time.
+with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called learning rate.
+Unlike supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time.
 This continuous evolution introduces instability during learning, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$.
 However, stability also depends on how the parameters $\theta$ change during each update.
-This work aims to adapt to these changes by dynamically selecting the learning rate $\eta$ over the training steps.
+This work aims to adapt to these changes by dynamically selecting the learning rate $\eta$ in the training steps.
 2.2 Multi-Armed Bandits
-MAB provide an elegant framework for making sequential decisions under uncertainty (see for instance 21).
+MAB provide an elegant framework for making sequential decisions under uncertainty (see for instance 22).
 MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$.
@@
 $G_{N}=\mathbb{E}\Biggl[\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)-\sum_{n=1}^{N}f_{n}(k_{n})\Biggr]\,$,
-where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)}$ represents the accumulated reward of the (pontentially changing) single best arm, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm.
-A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$: this guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
+where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)}$ represents the accumulated reward of the potentially changing single best arm, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm.
+A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$.
+This guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
 3 Dynamic Learning Rate for Deep RL
-In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic LRRL.
-LRRL is a meta-learning approach designed to dynamically select the learning rate in response to the agent’s performance.
-LRRL couples with SGD optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm.
-As the agent interacts with the environment, the average of observed rewards is used as bandit feedback to guide the selection of the most appropriate learning rate throughout the training process.
+In this section, we introduce LRRL, a meta-learning approach for dynamically adapting the learning rate based on policy performance.
+LRRL addresses the challenge of non-stationarity in RL by framing learning rate selection as a multi-armed bandit (MAB) problem.
+While the RL policy aims to maximize cumulative return, the bandit algorithm minimizes regret in identifying the most effective learning rates over time.
+Each candidate learning rate corresponds to an arm, and the bandit uses policy performance as feedback.
+We adopt an adversarial bandit algorithm with time-decay weighting to accommodate the evolving reward distribution as the agent interacts with the environment.
 3.1 Selecting the Learning Rate Dynamically
 Our problem can be framed as selecting a learning rate $\eta$ for policy updates —specifically, when updating the parameters $\theta$ after $\lambda$ interactions with the environment.
-Before training, a set $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ of $K$ learning rates are defined by the user.
+Before training, the user defines a set $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ of $K$ learning rates.
 Then, during training, a MAB algorithm selects, at every round $n$ —that is, at every $\kappa$ interactions with the environment— an arm $k_{n}\in\{1,\dots,K\}$ according to a probability distribution $p_{n}$ defined based on previous rewards, as explained in the next section.
@@
 Algorithm 1 dynamic Learning Rate for deep Reinforcement Learning (LRRL)
-Parameters:
-    Set of learning rates $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$
-    Number of episodes $M$
-    Horizon length $T$
-    Update window $\lambda$ for the learning network $\theta$
-    Update window $\tau$ for the target network $\theta^{-}$
-    Update window $\kappa$ for arm probabilities $p$
-Initialize:
-    Parameters $\theta$ and $\theta^{-}$
-    Arm probabilities $p_{0}\leftarrow(\tfrac{1}{K},\dots,\tfrac{1}{K})$
-    MAB round $n\leftarrow 0$
-    Cumulative reward $R\leftarrow 0$
-    Environment interactions counter $C\leftarrow 0$
+Initialize: network parameters $\theta$, target parameters $\theta^{-}$, arm probabilities $p_{0}\leftarrow(\tfrac{1}{K},\dots,\tfrac{1}{K})$, MAB round $n\leftarrow 0$, cumulative reward $R\leftarrow 0$, counter $C\leftarrow 0$
 for episode $m=1,2,3,\ldots,M$ do
   for timestep $t=1,2,3,\ldots,T$ do
-         Choose action $a_{t}$ following the policy $\pi_{\theta}(s)$ with probability $1-\epsilon$triangleright $(\epsilon$-greedy strategy)
+   Choose action $a_{t}$ following $\pi_{\theta}(\cdot|s)$ with probability $1-\epsilon$   $(\epsilon$-greedy strategy)
    Play action $a_{t}$ and observe reward $r_{t}$
    Add $r_{t}$ to cumulative reward $R\leftarrow R+r_{t}$
-         Increase environment interactions counter $C\leftarrow C+1$
+   Increase interactions counter $C\leftarrow C+1$
    if $C\mod\lambda\equiv 0$ then
      if $C\geq\kappa$ then
-                  Compute average of the last $C$ rewards $f_{n}\leftarrow\frac{R}{C}$
+      Compute average $f_{n}\leftarrow\frac{R}{C}$
       Increase MAB round $n\leftarrow n+1$
@@
       Sample arm $k_{n}$ with distribution $p_{n}$
-                  Reset $R\leftarrow 0$ and $C\leftarrow 0$
+      Reset $R\leftarrow 0$, $C\leftarrow 0$
      end if
@@
 end for
-Note that the same algorithm might be used with learning rates schedulers, that is with $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ where $\eta_{k}:\mathbb{N}\to\mathbb{R}_{+}$ is a predefined function, usually converging towards 00 at infinity.
+Note that the same algorithm might be used with learning rates schedulers, that is with $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ where $\eta_{k}:\mathbb{N}\to\mathbb{R}_{+}$ is a predefined function, usually converging towards 0 at infinity.
 If so, the learning rate used at round $n$ of the optimization is $\eta_{k_{n}}(n)$.
 3.2 Updating the Probability Distribution
-As we expect that the agent’s performance —and hence, the cumulative rewards— will improve over time, the MAB algorithm should receive non-stationary feedback.
-To take this non-stationary nature of the learning into account, we employ a modified version of the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction).
+Since the agent’s performance—and thus the cumulative reward—changes over time, the MAB algorithm naturally receives non-stationary feedback.
+To take this non-stationary nature of the learning into account, we employ a modified version of the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 5 for an introduction).
 At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively.
@@
 $f_{n}^{\prime}=f_{n}-\frac{1}{j}\sum_{i=0}^{j-1}f_{n-i}\,.$
-The improvement in performance allows computation of the next weights $w_{n+1}$ as follows, where initially $w_{1}=(0,\ldots,0)$:
-$\forall k\in\{1,\ldots,K\},\quad w_{n+1}(k)=\begin{cases}\delta\,w_{n}(k)+\alpha\,\frac{f^{\prime}_{n}}{e^{w_{n}(k)}}&\text{if }k=k_{n}\\ \delta\,w_{n}(k)&\text{otherwise}\,,\end{cases}$ (3)
-where $\alpha>0$ is a step-size parameter. The distribution $p_{n+1}$, used to draw the next arm $k_{n+1}$, is
+The improvement in performance allows the computation of the next weights $w_{n+1}$ as follows:
+[…]
+where initially $w_{1}=(0,\ldots,0)$ and $\alpha>0$ is a step-size parameter. The distribution $p_{n+1}$, used to draw the next arm $k_{n+1}$, is defined by:
 $\forall k\in\{1,\ldots,K\},\quad p_{n+1}(k)=\frac{e^{w_{n+1}(k)}}{\sum_{k^{\prime}=1}^{K}e^{w_{n+1}(k^{\prime})}}\,.$ (4)
@@
 4 Experiments
-In the following sections, we examine whether combining LRRL with Adam or RMSProp —two widely used optimization algorithms— can enhance cumulative returns in deep RL algorithms.
-We compare LRRL against learning methods both with and without schedulers, using baseline algorithms provided in the Dopamine [9] framework.
-The experiments involve testing LRRL across various configurations and a selection of Atari games, reporting the mean and one-half standard deviation of returns over 5 independent runs.
-As testing the full Atari benchmark is extremely costly, we choose a representative subset of games guided by their baseline learning curves, which offer insights into reward sparsity, stochasticity, and exploration demands.
-Further details on the evaluation metrics and hyperparameters used in these experiments are provided in Appendices C and D.
-In addition, although our primary target is RL tasks, we also test LRRL using SGD to minimize stationary objective functions to provide a simpler evaluation, as performance in RL is highly multi-factorial.
-4.1 Comparing LRRL with Standard Learning
-In our first experiment, we consider a set of $5$ learning rates.
-We set different configurations of LRRL using these values and compare them against the DQN agent reaching the best performance, in terms of maximum average return, among the possible single rates.
-More specifically, the learning rates considered are
+In this section, we evaluate LRRL across both discrete and continuous control tasks, using multiple RL algorithms and optimizers.
+Our study includes Atari games (discrete control, DQN and IQN agents) with baselines provided by the Dopamine framework [10], and MuJoCo tasks (continuous control, PPO agents) with baselines implemented in TorchRL [8].
+For Atari, we report the mean and one-half standard deviation of returns over 5 independent seeds, while for MuJoCo we use 10 seeds per algorithm.
+Across all tasks, we compare LRRL against fixed learning rates and schedulers, combined with widely used optimizers such as Adam and RMSProp.
+As testing the full Atari benchmark is computationally prohibitive, we select a representative subset of games guided by their baseline learning curves, capturing a range of properties such as reward sparsity, stochasticity, and exploration difficulty.
+To further disentangle RL-specific confounders, such as environment stochasticity and exploration dynamics, we also evaluate LRRL with SGD on stationary non-convex optimization problems, providing a controlled setting where performance can be attributed solely to learning rate adaptation.
+Additional details on evaluation metrics and hyperparameters are provided in Appendices D and E.
+4.1 LRRL vs. Fixed Learning Rates: Robustness Across Candidate Sets
+In our first experiment, we evaluate whether LRRL can remain competitive with tuned baselines when learning rates are drawn from different candidate sets. We consider five possible learning rates,
 $\mathcal{K}(5)=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr\}\,.$
-The whole set $\mathcal{K}(5)$ of rates is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) values, and the last one ($\mathcal{K}_{\text{sparse}}(3)$) taking the lowest/middle/highest values.
-All experiments use the Adam optimizer, and we report the return based on the same number of environment iterations.
-The results in Figure 1 demonstrate that LRRL using the full set $\mathcal{K}(5)$ substantially outperforms the best-performing single learning rate in 2 out of 6 games (Seaquest and Breakout) while remaining competitive in the others.
-As shown in Figure 7 in Appendix B.1, the baseline DQN performance is highly sensitive to the learning rate choice, requiring extensive hyperparameter tuning to identify the optimal rate.
-In contrast, LRRL dynamically selects learning rates during training, avoiding the need for such tuning.
-Additionally, the $3$-arm bandit variants highlight that while LRRL can explore multiple learning rates within a single run, the specific set of rates available still impacts performance.
-Overall, these results illustrate the advantage of LRRL, which achieves similar or better performance than the best-tuned DQN baseline while eliminating the computational cost of manual learning rate tuning.
-Figure 1: Comparison between LRRL and standard DQN learning: 5 variants of LRRL with different learning rate sets (provided in Appendix D) are tested again the DQN algorithm reaching best performance among the possible learning rates.
-The mean and one-half standard deviations over 5 runs are respectively represented by the curves and shaded areas.
-Next, we illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves.
-Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$.
-In most of the tested environments, LRRL behaves similarly to time-decay schedulers by selecting higher learning rates during the early stages of training, gradually shifting toward arms with lower rates as training progresses.
+and construct multiple LRRL variants based on different subsets: the full set $\mathcal{K}(5)$, the three smallest rates $\mathcal{K}{\text{lowest}}(3)$, the three middle rates $\mathcal{K}{\text{middle}}(3)$, the three largest rates $\mathcal{K}{\text{highest}}(3)$, and a sparse mix of low, medium, and high values $\mathcal{K}{\text{sparse}}(3)$.
+All experiments use the Adam optimizer.
+Figure 1: LRRL and standard DQN comparison: 5 variants of LRRL with different learning rate sets are tested against the DQN algorithm reaching best performance among the possible learning rates (provided in Appendix 7).
+The mean and one-half standard deviations over 5 runs are represented.
+It is important to note that the performance of LRRL naturally depends on the composition of the candidate set.
+If most of the available learning rates perform poorly in isolation, then LRRL’s performance will also be limited, as it can only adapt within that set.
+This does not indicate a failure of the method, but rather reflects the quality of the available candidates.
+Conversely, when the set contains both effective and ineffective rates, LRRL consistently adapts toward the better-performing ones, maintaining stability even when poor candidates are present.
+In this experiment, we compare LRRL against the best single fixed rate among the candidates.
+This baseline represents an oracle that assumes prior knowledge of the best-performing learning rate, requiring multiple training runs.
+LRRL, by contrast, adapts within a single run.
+To better understand how this adaptation occurs, we analyze the sequence of learning rates selected during training.
+Figure 2 visualizes the arms pulled by LRRL $\mathcal{K}(5)$ and the corresponding returns from a single run.
+In most environments, LRRL initially favors higher learning rates to accelerate early progress, before gradually shifting toward smaller ones as performance stabilizes.
+This behavior resembles a decay schedule, but crucially, it emerges from policy performance feedback rather than being predefined.
+The ability to combine aggressive and conservative updates within a single trajectory illustrates how LRRL dynamically adapts to non-stationary learning dynamics.
 Figure 2: Systematic sampling of normalized learning rates and returns over the training steps using LRRL $\mathcal{K}(5)$ with Adam optimizer, through a single run.
@@
 Lower rates are increasingly selected over time.
-4.2 Combining and Comparing Schedulers with LRRL
-In the following experiments, LRRL is combined with learning rate schedulers.
-Specifically, we employ schedulers with exponential decay rate of the form $\eta(n)=\eta_{0}\times e^{-d\,n}$, where $\eta_{0}$ is a fixed initial value (common to each scheduler and equal to $6.25\times 10^{-5}$ in our experiment), $d$ is the exponential decay rate and $n$ is the number of policy updates (i.e., of MAB rounds).
-We define a set of $3$ schedulers $\mathcal{K}_{s}$, where each arm represents a scheduler using a different decay rate $d=\{1,2,3\}\times 10^{-7}$, and compare the results of LRRL with each scheduler individually, using the Adam optimizer.
-Figure 3 shows that LRRL combined with schedulers could once again substantially increase the final performance of 2 out of 6 tasks compared to using a fixed scheduler.
-In the remaining tasks, LRRL remains at least competitive, while for most of them, it shows a better jumpstart performance.
-The dashed black line represents the maximal average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by Andrychowicz et al.
-[4], who linearly decay the learning rate to 00.
-In all the experiments conducted so far, we have set the update window $\kappa$ parameter to 1, which allows the observation of a full episode before updating the learning rate.
-In Appendix B.3, we address how this value could influence the performance of LRRL by allowing the observation of more interactions.
-Figure 3: LRRL with arms that select learning rate schedulers: LRRL vs 3 individual schedulers ($d=1,2,3$).
-The dashed black line represents the max average return achieved by Adam with a constant learning rate set to $6.25\times 10^{-5}$.
-4.3 Comparing RMSProp and Adam Optimizers
-Another widely used optimizer for deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism.
-Adam builds upon RMSProp by retaining exponential moving averages to give more weight to recent gradients while incorporating momentum.
-Although the standard RMSProp does not feature momentum, we found that adding momentum to RMSProp can increase both the performance of DQN and LRRL, aligning with findings in the literature [30, 4].
-In the following experiment, we compare the performance of RMSProp with Nesterov’s momentum (RMSProp-M), and Adam when coupled with either LRRL or the best-performing single learning rate when using DQN.
-Since the results were consistent in previous sections, we extend the experiments to other 3 Atari games.
-Figure 4 shows that LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN.
-Moreover, LRRL (RMSProp-M) underperforms compared to DQN without LRRL in two out of three tasks due to its slow convergence despite better jumpstart performance.
-Future work should investigate whether this slow convergence is due to some optimizer features such as the absence of bias correction in the first and second moment estimates.
-Figure 4: A comparison between Adam and RMSProp with momentum, using either DQN or LRRL.
-4.4 LRRL with a Distributional RL Algorithm
-Lastly, we illustrate the versatility of LRRL by combining it with the distributional RL algorithm Implicit Quantile Networks (IQN) [11], which exhibits the best performance in many Atari games among the algorithms available in Dopamine.
-Adam is used as the optimizer.
-Figure 5 shows that LRRL outperforms standard learning with a single learning rate across all the tasks evaluated.
-More generally, LRRL can be seamlessly integrated with any RL algorithm that employs function approximators based on SGD, improving its performance and mitigating hyperparameter tuning.
-Figure 5: Comparison between LRRL and standard IQN learning: IQN combined with LRRL using 3 different learning rates is tested against the 3 variants of standard IQN using each possible learning rate.
+4.2 LRRL with Learning Rate Schedulers
+We next investigate whether LRRL can improve upon fixed learning rate schedulers by adapting among them during training.
+Specifically, we consider exponential decay schedules of the form $\eta(n)=\eta_{0}\times e^{-dn}$, where $\eta_{0}=6.25\times 10^{-5}$ is a fixed initial value, $d$ is the exponential decay rate, and $n$ is the number of policy updates (i.e., MAB rounds).
+We define a candidate set $\mathcal{K}_{s}$ of three schedulers with $d\in{1,2,3}\times 10^{-7}$, and compare LRRL against each scheduler used individually, with Adam as the optimizer.
+Figure 3 shows that LRRL achieves clear gains in two out of six tasks and remains competitive in the others.
+Importantly, LRRL consistently improves jumpstart performance in most of tasks.
+This highlights a key advantage over fixed schedulers: whereas decay schedules impose a monotonic reduction in learning rate, LRRL adapts the choice dynamically based on policy performance.
+For reference, the dashed black line indicates the performance of Adam with a constant learning rate, which is generally slightly weaker than the best-performing decayed scheduler, in line with previous findings [3].
+Figure 3: LRRL vs 3 individual schedulers ($d=1,2,3$). The dashed black line represents the max average return achieved by Adam with a constant learning rate set to $6.25\times 10^{-5}$.
+Effect of the Update Window $\kappa$.
+In the experiments so far, we have fixed the update window $\kappa$ to $1$, meaning LRRL observes a complete episode before updating the learning rate.
+However, in highly stochastic environments, updating after every episode may provide noisy feedback, making it harder for LRRL to identify the most effective learning rate.
+Increasing $\kappa$ allows more interactions before each update, trading off reactivity for stability.
+Two factors should guide this choice:
+  • Network update rate: If learning rates are switched more frequently than model updates occur, the feedback to the bandit become inaccurate.
+  • Environment feedback: In tasks where outcomes are only revealed after long action sequences, observing more steps before updating helps capture agent performance more reliably.
+We evaluate LRRL with different values of $\kappa$ when combined with schedulers.
+As shown in Figure 4, the effect of $\kappa$ is task-dependent: some games benefit from more reactive updates, while others gain from longer observation windows.
+Overall, $\kappa=1$ remains a robust choice for Atari games, though higher values can yield slight improvements in certain tasks.
+Figure 4: Results of the ablation study evaluating the impact of varying $\kappa$ with LRRL using schedulers.
+The dashed black line represents the max average return achieved by the best performing scheduler.
+4.3 Continuous Action Spaces
+In the following, we evaluate LRRL on MuJoCo tasks using Proximal Policy Optimization (PPO) [36] implemented in TorchRL [8], which by default applies a linear decay schedule to both actor and critic learning rates.
+We compare LRRL against two alternative strategies: AdamRel [15] and Cyclical Learning Rates (CLR) (see details in B.5), the latter of which has shown promising results in RL despite higher instability [26].
+To isolate the effect of the adaptation mechanism (rather than LR range), we adopt the following implementation choices for this comparison: (i) CLR cycles deterministically over the same candidate learning rate set used by LRRL; (ii) AdamRel uses the standard fixed learning rate (no decay) in this experiment [^1] AdamRel could be combined with decay; CLR could cycle over other ranges.
+We fix these choices to focus on the adaptation criterion.; and (iii) actor and critic share the same learning rate across methods.
+We report means and one-half standard deviations over 10 seeds, evaluating policies after each batch for a total of 1M frames.
+As shown in Figure 5, LRRL achieves higher returns on two of five MuJoCo tasks.
+Unlike CLR, which cycles through rates without a performance-based criterion, LRRL adapts its selection from the same pool using policy feedback, leading to stronger or comparable outcomes in most of the tasks.
+Compared to AdamRel, LRRL does not rely on optimizer-specific mechanics (momentum resets) and is applicable across different optimizers while still adapting to non-stationary learning dynamics.
+Figure 5: Comparison of LRRL, AdamRel, and CLR on MuJoCo tasks using PPO.
+LRRL achieves higher average returns in two tasks.
+Lines represent the mean and shaded regions show one-half standard deviations over 10 seeds, with training run for 1M frames.
 4.4 SGD and Stationary Non-convex Loss Landscapes
-Although RL tasks are a challenging showcase to illustrate the effectiveness of our approach in the non-stationary case, there are many confounding factors to analyze their impact on the results, such as the optimizer adaptive routine, the environment stochasticity, and the reward sparseness.
-To disentangle these factors from our evaluation, in the following experiments we employ the SGD optimizer, which has a single scalar learning rate without adaptive scheme or momentum to find the global minima of $6$ non-convex loss landscapes.
-Since these functions represent a stationary case, we employ the stochastic bandit Minimax Optimal Strategy in the Stochastic case (MOSS) algorithm (see details in Appendix B.2), using as feedback the loss plus a small constant $\xi$, such that $f_{n}=\frac{1}{|\textrm{loss}(n)|+\xi}$, observed at each gradient step $n$ to update the arm’s probabilities.
-Figure 6 illustrates the performance of LRRL against SGD with a single learning rate, starting from the same initial point and for the same amount of steps.
-The final steps are indicated by a star icon with a different color.
-The results show that LRRL, with its “all-in-one” strategy, efficiently navigates intricate landscapes.
-It is effective in combining higher learning rates to accelerate convergence and lower ones for more fine-grained optimization steps when needed, as evidenced in the loss landscapes depicted in Figures 6(a)6(b), and 6(c).
-In addition, Figures 6(d)6(e), and 6(f) illustrate that LRRL can be also effective in converging towards the best-performing learning rate within the arms, showcased by its steady convergence towards the global minima.
-Figure 6: Comparison between LRRL and fixed learning rates in a stationary setting with SGD.
-LRRL with 3 learning rates is tested against the 3 constant rates for 6 non-convex losses (shown in Figure 10 in Appendix B.4).
+Although RL tasks are a natural showcase for LRRL under non-stationarity, they involve many confounding factors, such as the adaptive routine of the optimizer, stochastic environment dynamics, and reward sparsity.
+To disentangle these effects, we evaluate LRRL in a simpler, stationary setting: minimizing six standard non-convex loss landscapes using stochastic gradient descent (SGD).
+Unlike Adam or RMSProp, SGD employs a single scalar learning rate without momentum or adaptivity, making it an ideal testbed to isolate the impact of dynamic learning rate selection.
+Experimental setup.
+Each minimization works for a fixed budget of steps from the same initialization.
+LRRL is applied on top of SGD using the stochastic bandit Minimax Optimal Strategy in the Stochastic case (MOSS; see Appendix B.2), with arms corresponding to candidate learning rates.
+At each gradient step $n$, the bandit receives the feedback
+$f_{n}=\frac{1}{\mathcal{J}(w_{1}^{(n)},w_{2}^{(n)})+\xi},\quad\xi>0,$
+where $\mathcal{J}$ denotes the loss and $\xi$ prevents division by zero.
+The bandit updates after every gradient step.
+Figure 6 illustrates the optimization trajectories, with final iterates marked by stars.
+Two consistent behaviors emerge:
+  • Adapting across multiple rates.
+On landscapes such as Beale, Bohachevsky, and Griewank, LRRL flexibly alternates between larger and smaller learning rates depending on the stage of optimization.
+By “playing” with multiple rates, LRRL can accelerate progress when higher values are beneficial and switch to smaller ones when finer convergence is required.
+This adaptability allows it to escape poor local minima and achieve stronger convergence than any single fixed rate.
+  • Converging to the best candidate.
+On functions where a single rate is clearly optimal (e.g., Rosenbrock, Zakharov), LRRL adapts toward that rate, outperforming the second best performing candidate.
+These experiments demonstrate that LRRL is not only robust to candidate quality but can also exploit multiple learning rates within a single run.
+Whereas a fixed learning rate must trade off between speed of convergence and stability, LRRL adapts dynamically, combining both advantages.
+Importantly, these results in stationary optimization confirm that LRRL’s effectiveness is not limited to the non-stationary structure of RL, but reflects a more general capacity to adapt learning dynamics.
+Figure 6: LRRL vs fixed learning rates in a stationary setting with SGD.
+LRRL with 3 rates is tested against the 3 constant rates for 6 non-convex losses (shown in Figure 11 in Appendix).
 LRRL is effective in either converging to the best performing rate or combining the set to achieve better results.
 5 Conclusion
-In this work, we introduced LRRL, a meta-learning approach for selecting optimizers’ learning rates on the fly.
-In any optimization task, the best-performing learning rate or rate scheduler is typically unknown, requiring multiple runs to test each candidate in a wasteful computational effort.
-LRRL mitigates this cost by employing a multi-armed bandit algorithm to explore a set of learning rates within a single run, significantly reducing the need for extensive hyperparameter tuning while achieving better or competitive results with minimal computational overhead.
-Using RL and non-convex optimization tasks, our empirical results show that combining LRRL with the Adam or SGD optimizers could enhance the performance, outperforming baselines and learning rate schedulers.
-Future investigations could extend LRRL ideas to other hyperparameters, such as mini-batch size, which also plays a key role in convergence.
-Moreover, although LRRL selects rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [33, 19, 1].
+We introduced LRRL, a meta-learning approach that adapts the optimizer’s learning rate on the fly using a multi-armed bandit.
+In contrast to fixed schedules or oracle tuning that require multiple runs, LRRL evaluates and selects among candidate rates within a single run, reducing tuning effort while remaining competitive or superior to the best fixed choice.
+Across discrete and continuous RL tasks, LRRL improved or matched strong baselines such as decay schedules, AdamRel, and CLR, and generalized to stationary non-convex optimization with SGD.
+These results highlight LRRL’s ability to combine aggressive and conservative updates within a single trajectory, offering both robustness and adaptability.
 Limitations.
-The primary limitation of LRRL lies in the need to define a set of arms for the MAB algorithm, as the choice of this set can directly influence the performance of the underlying learning algorithm.
-However, in computationally intensive settings such as RL tasks, LRRL offers a significant advantage: it enables the simultaneous evaluation of multiple learning rates within a single training run, eliminating the need to test each rate individually.
-By employing an “all-in-one” strategy, LRRL not only achieves competitive results while reducing the burden of hyperparameter optimization but also can outperform standard learning by combining multiple learning rates.
+LRRL operates over a predefined set of candidate learning rates, which serves as the search space for adaptation.
+However, this requirement is inherent to any hyperparameter search, and our experiments show that LRRL can exploit multiple candidates dynamically, avoiding catastrophic failures and often outperforming fixed-rate baselines.
+Theoretical guarantees would require strong assumptions (e.g., convexity, smoothness, bounded losses) that deviate from deep RL; instead, we focused on broad empirical validation.
 Broader Impact Statement
@@
 This section presents works closely related to ours.
-We first introduce approaches that use MAB for hyperparameter selection and later other methods that aim to adapt the learning rate over the training process.
+We first introduce approaches that use MAB for hyperparameter selection, and later other methods that aim to adapt the learning rate over the training process.
 Multi-armed bandit for (hyper)parameter selection.
@@
 Moreover, it also introduces an exploration strategy which depends on epistemic uncertainty (captured through ensembles) and aleatoric uncertainty (estimated using a distributional RL algorithm).
-Second, TOP trades-off between optimism and pessimism by adding new variables (arms), which are not inherently tied to the learning process and also require tuning.
+Second, TOP trades-off optimism and pessimism by adding new variables (arms), which are not inherently tied to the learning process and also require tuning.
 In contrast, LRRL offers an algorithm-agnostic solution by directly optimizing the learning rate, a core hyperparameter in all SGD-based methods, making it simpler and more broadly applicable.
@@
 [24] propose Adam with Bandit Sampling (ADAMBS), which employs the Exp3 algorithm to enhance sample efficiency by incorporating importance sampling within the Adam optimizer.
-ADAMBS prioritizes informative samples by keeping a probability distribution over all samples, although the feedback is provided by the gradient computed over a mini-batch, unlike previous works in RL that are capable of assigning per-sample importance by using their computed TD-error (see 34).
+ADAMBS prioritizes informative samples by keeping a probability distribution over all samples, although the feedback is provided by the gradient computed over a mini-batch, unlike previous works in RL that are capable of assigning per-sample importance by using their computed TD-error (see 35).
 In order to select the learning rate for stochastic gradient MCMC, Coullon et al.
-[10] employ an algorithm based on Successive Halving [17, 16], a MAB strategy that promotes promising arms and prunes suboptimal ones over time.
+[11] employs an algorithm based on Successive Halving [19, 18], a MAB strategy that promotes promising arms and prunes suboptimal ones over time.
 In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [23, 38].
 A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate on the fly from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space.
-Close to our work, Parker-Holder et al.
-[29] propose population-based training which employs a Gaussian Process (see details in 31) with a time-varying kernel to represent the upper confidence bound in a bandit setting.
-The idea is to have many instances of agents learning a task in parallel with different hyperparameters, replacing those that are underperforming within a single run.
+Parker-Holder et al.
+[32] propose population-based training which employs a Gaussian Process (see details in 34) with a time-varying kernel to represent the upper confidence bound in a bandit setting.
+The idea is to have multiple instances of agents learning a task in parallel, each with different hyperparameters, and replace those that are underperforming within a single run.
+Close to our work, MABSearch [43] frames the learning rate selection as a MAB.
+In practice, however, its algorithm reduces to an heuristic $\epsilon$-greedy scheme that maintains an exponential moving average of the loss to guide exploitation.
+As such, its performance depends critically on the choice of $\epsilon$ and its decay, lacking the principled, adaptive exploration–exploitation balance over the long run that defines proper bandit methods.
+In contrast, LRRL builds on adversarial algorithms such as Exp3, which maintain probabilistic weights over arms and update them under non-stationary feedback, providing the first rigorous integration of bandits as a meta-optimizer for learning rate adaptation.
 Learning rate adapters/schedulers.
 Optimizers such as RMSProp [45] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [20] also incorporates momentum to smooth gradient steps.
-However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see 13).
+However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see 14).
 Increment-Delta-Bar-Delta (IDBD), introduced by Sutton [41], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [51].
@@
 A meta-gradient RL method is proposed in Xu et al.
-[45], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function.
-Our work differs from these methods by employing a MAB approach to dynamically select the learning rate over the training process, specifically targeting RL settings.
+[49], formulated as a two-level optimization process: one level optimizes the agent’s policy objective, while the other learns meta-parameters of the objective function.
+Although both LRRL and meta-gradient approaches aim to enable agents to “learn how to learn” within a single agent’s lifetime, meta-gradient methods must be tailored to the specific loss structure and update rules of each RL algorithm.
+In contrast, LRRL is algorithm-agnostic: it operates at the optimizer level, adapting the learning rate through a bandit mechanism without requiring gradients of the learning objective or modifications according to the underlying algorithm.
 Appendix B Supplementary Experiments
 B.1 Baseline Evaluation with Varying Learning Rates
-To establish a baseline to compare learning without our approach LRRL, we run individual arm values as baseline learning rate using the Adam optimizer.
-The results presented in Figure 7 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to worse jumpstart performance and slow convergence.
-Figure 7: DQN performance using the Adam optimizer with varying learning rates across 6 Atari games.
-Shaded region represents one-half standard deviation.
-The results show that the best performing learning rate varies across games.
-B.2 A Comparison between Multi-Armed Bandits Algorithms
-Adversarial MAB algorithms are designed for environments where the reward distribution changes over time.
-In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown probability distributions.
-To validate our choice and address our method’s robustness, we compare Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [5].
-MOSS trades off exploration and exploitation by pulling the arm $k$ with highest upper confidence bound given by:
+To establish baselines, we evaluate DQN with Adam across several Atari tasks using fixed learning rates.
+As shown in Figure 7, larger learning rates may provide a better jumpstart but often lead to worse asymptotic performance, while smaller rates result in more stable convergence but slower learning, with the best-performing learning rate varying by environment.
+This illustrates why a dynamic strategy such as LRRL, which is able to exploit higher rates early and smaller ones later, can be advantageous.
+Figure 7: Baseline evaluation of DQN with Adam using different fixed learning rates across Atari tasks.
+Performance varies substantially across games, confirming that no single learning rate works universally.
+B.2 Comparing adversarial and stochastic MAB algorithms
+Adversarial MAB algorithms are designed for settings where reward distributions can change over time, making them naturally suited to RL.
+In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown distributions.
+To validate our design choice and assess robustness, we compare LRRL using Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [4].
+MOSS balances exploration and exploitation by selecting the arm $k$ with the largest upper confidence bound:
 $B_{k}(n)=\hat{\mu}_{k}(n)+\rho\,\sqrt{\frac{\max\Bigl(\log{\frac{n}{Kn_{k}(n)}},0\Bigr)}{n_{k}(n)}}$
-where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to round $n$.
-In Figure 8, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balances exploration and exploitation in MOSS.
-Additionally, the bandit feedback used in MOSS is the average cumulative reward $\frac{R}{C}$.
-The results indicate that while Exp3 performs better overall, MOSS can still achieve competitive results depending on the amount of exploration, demonstrating the robustness of our approach regarding the MAB algorithm employed.
-Figure 8: A comparison between adversarial (Exp3) and stochastic (MOSS) MAB algorithms. Shaded region represents one-half standard deviation. The results show that Exp3 performs better overall.
-B.3 Dealing with Noise and Lack of Feedback
-LRRL relies entirely on performance improvements to update the arm probabilities.
-However, some RL environments can be highly stochastic, with sparse or delayed rewards, which increases the difficulty of learning tasks and, consequently, the ability of LRRL to make optimal choices.
-To mitigate this limitation, we introduce the variable $\kappa$ in Algorithm 1, which determines the number of interactions to observe before making a decision.
-When setting $\kappa$, two factors should be considered:
-  • The network’s update rate: Selecting new learning rates more frequently than updating the model leads to inaccurate feedback for the bandit mechanism since no model update might occur after pulling some arm.
-  • The number of environment steps: In some tasks, positive or negative outcomes of a sequence of actions are only perceived at the end of an episode, and capturing sufficient information about the agent performance is essential to updating the arms’ probabilities.
-In the experiments presented in this work, we set $\kappa=1$ episode, allowing the observation of a complete episode before updating the arms’ probabilities.
-To explore the impact of this parameter further, we conduct an ablation study using different $\kappa$ values to evaluate potential performance improvements when combined with the schedulers shown in Figure 3.
-As suggested by Figure 9, the choice of $\kappa$ is task-dependent: while some tasks benefit from more reactive learning rate updates, others show marginal improvements with increased observation counts.
-Ultimately, setting $\kappa=1$ seems relatively robust for Atari games, although higher values could slightly increase the performance in some tasks.
-Figure 9: Results of the ablation study evaluating the impact of varying $\kappa$ with LRRL using schedulers.
-The dashed black line represents the max average return achieved by the best performing scheduler.
-B.4 Supplementary Results using SGD
-In Section 4.5, we tested the SGD optimizer in $6$ non-convex loss landscapes show in Figure 10.
-These functions were selected since they represent common benchmarks in optimization, each one offering different challenges due to their distinct shapes.
-By using SGD on these landscapes, we minimize the influence of confounding factors, such as the high stochasticity inherent in some RL tasks and the complexity of more robust optimizers like Adam.
-Figure 11 presents the normalized loss and the arms selected at each optimization step.
-The results demonstrate LRRL’s dynamic adaptation mechanism, either converging to the most effective learning rate or combining them to achieve superior outcomes.
-Figure 10: The 6 stationary non-convex loss landscapes used to assess LRRL with SGD optimizer.
-Figure 11: Learning rates and logarithmic scale of losses over the gradient steps using LRRL with SGD optimizer.
-Appendix C Evaluation Metrics
+where $\hat{\mu}_{k}(n)$ is the empirical average reward of arm $k$, and $n_{k}(n)$ is the number of times it has been pulled up to round $n$.
+In our experiments (Figure 8), we vary the step-size $\alpha$ for Exp3 and the exploration parameter $\rho$ for MOSS, using average cumulative return $f_{n}$ as feedback for MOSS.
+Results show that Exp3 achieves stronger overall performance, consistent with the non-stationary nature of RL.
+Nevertheless, MOSS can reach competitive outcomes under appropriate exploration settings, demonstrating that LRRL is robust to the choice of bandit algorithm.
+Figure 8: Comparison of adversarial (Exp3) and stochastic (MOSS) bandit algorithms within LRRL.
+Exp3 performs better overall, as expected in non-stationary RL settings, but MOSS can achieve competitive results under certain exploration parameters, highlighting robustness to the bandit choice.
+B.3 Comparing RMSProp and Adam Optimizers
+We next evaluate LRRL’s robustness to the choice of optimizer.
+We compare Adam and RMSProp augmented with Nesterov momentum (RMSProp-M), since prior work [3] and our own observations indicate that adding momentum consistently improves RMSProp’s performance.
+Both Adam and RMSProp-M are paired with either LRRL and the best-performing fixed learning rate within the DQN agent.
+Since LRRL has already shown consistency in earlier sections, we extend the evaluation to three additional Atari games here.
+As shown in Figure 9, LRRL combined with Adam achieves the strongest results overall, outperforming both RMSProp-M baselines and the tuned fixed-rate Adam baseline.
+LRRL with RMSProp-M provides better jumpstart but converges more slowly, underperforming standard DQN in two out of the three tasks.
+One possible explanation is the lack of bias correction in RMSProp’s moment estimates, which may limit LRRL’s ability to fully adapt.
+Exploring such optimizer-specific interactions is an interesting direction for future work.
+Figure 9: Comparison of Adam and RMSProp with Nesterov momentum (RMSProp-M), each combined with either LRRL or the best-performing fixed learning rate on DQN across three additional Atari games.
+LRRL with Adam consistently achieves the strongest results, while LRRL with RMSProp-M shows good jumpstart performance but slower convergence.
+B.4 Distributional RL: Implicit Quantile Networks (IQN)
+To further illustrate the generality of LRRL, we combine it with Implicit Quantile Networks (IQN) [12], a distributional RL algorithm available in Dopamine that has shown stronger empirical performance than DQN and other algorithms in several Atari tasks.
+Figure 10 shows that LRRL improves or matches performance across the evaluated Atari tasks when paired with IQN.
+These results demonstrate that LRRL is not tied to a specific algorithmic family: beyond standard value-based methods such as DQN or actor–critic methods such as PPO, LRRL can also enhance distributional agents.
+More broadly, since LRRL operates at the optimizer level by dynamically selecting the learning rate, it can in principle be coupled with any algorithm based on SGD or its variants.
+This makes LRRL widely applicable across deep RL.
+Figure 10: IQN with Adam on Atari tasks: LRRL vs.
+fixed learning rates.
+LRRL improves or matches performance across environments, showing that dynamic learning-rate adaptation benefits stronger distributional RL agents as well as DQN.
+B.5 Supplementary Results using SGD
+Figure 11: The six stationary non-convex loss landscapes used to assess LRRL with the SGD optimizer.
+Each function presents distinct challenges, such as narrow valleys (Rosenbrock) or multiple local minima (Griewank).
+(a) Beale
+(b) Rosenbrock
+(c) Three-Hump Camel
+(d) Zakharov
+Figure 12: Convergence steps toward the global minimum using LRRL and CLR on four benchmark functions.
+In Section 4.4, we tested LRRL with SGD on six non-convex loss landscapes (Figure 11).
+These benchmarks are widely used in optimization because their diverse shapes highlight different challenges, such as multiple local minima, steep ridges, or flat plateaus.
+Using SGD in this stationary setting minimizes the influence of confounders such as environment stochasticity in RL or the adaptivity of more sophisticated optimizers like Adam.
+Here, we extend our experiments and compare LRRL against Cyclical Learning Rates (CLR) [40].
+CLR is motivated by a similar intuition that varying learning rates can improve performance, but it follows a predetermined periodic schedule that cycles through candidate values with a fixed step size $n$ (see Algorithm 2).
+Unlike LRRL, CLR does not use feedback to guide its choices.
+As a result, while CLR can occasionally reach strong solutions, it may introduce instability by switching rates at arbitrary times.
+In contrast, LRRL adapts its selection directly from performance feedback.
+Algorithm 2 Cyclical Learning Rates [40]
+1:Input:
+2:$n$: step size (half-cycle length in iterations)
+3:minlr: minimum learning rate
+4:maxlr: maximum learning rate
+5:$M$: total number of training steps
+6:for $m=1,2,\dots,M$ do
+7:  $\text{cycle}\leftarrow\left\lfloor 1+\frac{m}{2n}\right\rfloor$   (Compute current half-cycle index)
+8:  $x\leftarrow\left|\frac{m}{n}-2\cdot\text{cycle}+1\right|$   (Normalized position within cycle)
+9:  $\eta\leftarrow\text{minlr}+(\text{maxlr}-\text{minlr})\cdot\max(0,1-x)$   (Compute learning rate)
+10:end for
+Figure 12 shows the gradient steps taken employing the best performing candidate, in terms of convergence rate, varying respectively the exploration trade-off given by $\rho$ (LRRL) and the step granularity $n$ (CLR).
+Table 1 displays the convergence steps to global minima considering loss landscapes where that could be achieved with an Euclidean distance less than $0.01$.
+Notably, LRRL has superior outcomes despite its higher sensitivity to the hyperparameter choice.
+Loss Landscape CLR ($n=2$) CLR ($n=5$) CLR ($n=10$) LRRL ($\rho=1$) LRRL ($\rho=2$) LRRL ($\rho=3$)
+Beale - 1268 1248 2083 905 907
+Rosenbrock 96267 96267 96275 65318 118167 66257
+Three-Hump Camel 292 320 359 286 518 169
+Zakharov 237 237 236 171 284 178
+Table 1: Convergence steps to global minima (Euclidean distance <0.01<0.01) for different loss landscapes and learning rate schedules.
+Appendix C Additional Discussion and Limitations
+C.1 On Selecting the Learning Rates and Hyperparameters
+Learning rates.
+Figure 7 shows that different learning rates lead to very different policy performance, yet there is no principled way to select them without extensive testing.
+LRRL does not eliminate the need to define a candidate set, and its performance naturally depends on this choice, as seen in Figure 1.
+However, LRRL reduces the burden of tuning by adaptively selecting among candidates within a single run, instead of requiring multiple runs to evaluate each rate individually.
+As illustrated in the SGD experiments (Figure 6), this allows LRRL to combine aggressive and conservative rates dynamically or to converge toward the best rate when one dominates.
+Hyperparameters.
+For the bandit algorithm, we fixed $\delta=0.9$ in all experiments, a common choice for decay factors (e.g., momentum or the RL discount factor $\gamma$).
+For the bandit feedback $f^{\prime}_{n}$, a history window of $j=5$ was robust across environments.
+As shown in Figure 4, $\kappa$ only slightly affects performance, while $\alpha$ is the most sensitive parameter (Figure 8), requiring tuning.
+This reflects a broader challenge in meta-optimization: hyperparameters are themselves required to control the process of tuning other hyperparameters [27].
+C.2 The Cumulative Reward as a Proxy of Performance
+Figure 13: Return and loss in RL.
+A central design choice in LRRL is the feedback used by the bandit algorithm.
+Our setting corresponds to a dual objective: the bandit aims to minimize regret in selecting learning rates, while the RL agent seeks to maximize cumulative returns.
+We therefore use the improvement in policy performance $f^{\prime}_{n}$ as feedback to update arm probabilities.
+Unlike supervised learning, where the training loss directly reflects predictive quality, in RL the optimization loss is only loosely related to policy performance (Figure 13).
+For example, the TD-error in Equation 1 measures how surprising a transition is, but a high TD-error does not necessarily indicate that the policy is improving.
+By contrast, returns directly measure policy effectiveness, even if they are noisy or sparse.
+This noise is intrinsic to RL and affects both LRRL’s feedback and policy optimization itself.
+C.3 Computational Overhead
+The overhead introduced by LRRL is governed primarily by the decision frequency $\kappa$ (how often we update arm probabilities) and the number of arms $K$.
+Each bandit update costs $O(K)$ for computing/sampling the distribution, so the total cost over $T$ decisions is $O(TK)$.
+In our SGD experiments, we update once per gradient step ($\kappa=1$ step), so the extra $O(K)$ work can be noticeable when the model is tiny (e.g., 2D landscapes) because the gradient step itself is very cheap.
+In RL, we set $\kappa=1$ episode while policy updates happen every $\lambda=4$ steps with mini-batches; thus $T$ is roughly the number of episodes.
+With small $K$ (as used here), the $O(TK)$ bandit cost is negligible compared to forward/backward passes and replay/batching, and adds no meaningful wall-clock overhead.
+C.4 Guarantees
+While adversarial bandit algorithms offer theoretical regret guarantees under certain conditions, these guarantees do not carry over to deep RL because the best-performing arm itself changes as the policy evolves.
+From a theoretical point of view, guarantees would require assumptions (e.g, convexity, smoothness, bounded losses) that deviate significantly from the realities of RL.
+We therefore focus on empirical validation across diverse benchmarks, demonstrating LRRL’s robustness and generality in practice.
+C.5 Future Work
+A key limitation of LRRL is the need to predefine a set of candidate learning rates, although our experiments show some robustness to this choice.
+One possible direction to overcome this limitation is using Bayesian optimization, for example with Gaussian Processes (GPs) [34].
+While GPs are sample-efficient, their computational cost increases as the number of observations grows, and they can struggle with non-stationary objectives.
+Addressing this would likely require mechanisms that emphasize recent interactions, as in time-varying GP-UCB [7], or approaches that limit and/or adapt the number of inducing points [33].
+Integrating such techniques with LRRL could reduce reliance on a fixed candidate set while retaining adaptability to non-stationary dynamics.
+Appendix D Evaluation Metrics
 In this section, we describe the evaluation metrics that can be used to evaluate the agent’s performance as it interacts with an environment.
 Based on Dopamine, we use the evaluation step-size “iterations”, defined as a predetermined number of episodes.
+In MuJoCo tasks, updates refer to policy updates.
 Figure 14 illustrates the evaluation metrics used in this work, as defined in Taylor and Stone [44]:
@@
 A lower jumpstart performance can result from factors such as a lower learning rate, although this work demonstrates that this does not necessarily lead to worse final performance.
-Figure 12: Performance curves of two RL algorithms (adapted from 40).
-Appendix D Implementation Details
+Figure 14: Performance curves of two RL algorithms (adapted from 44).
+Appendix E Implementation Details
 In the following, we list the set of arms, optimizer and the bandit step-size used in each experiment.
-Section 4.1 – Comparing LRRL with Standard Learning.
+Section 4.1 – LRRL vs. Fixed Learning Rates: Robustness Across Candidate Sets
   • Optimizer: Adam
@@
   • Considered sets of learning rates:
-    $\mathcal{K}(5)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
-    $\mathcal{K}_{\text{lowest}}(3)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
-    $\mathcal{K}_{\text{middle}}(3)$ =$=\hskip 65.44142pt\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
-    $\mathcal{K}_{\text{highest}}(3)$ =$=\hskip 125.19194pt\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
+    $\mathcal{K}(5)$ $=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr\}$
+    $\mathcal{K}_{\text{lowest}}(3)$ $=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr\}$
+    $\mathcal{K}_{\text{middle}}(3)$ $=\hskip 65.44142pt\bigl\{3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr\}$
+    $\mathcal{K}_{\text{highest}}(3)$ $=\hskip 125.19194pt\bigl\{6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr\}$
 […]
-Section 4.2 – Combining and Comparing Schedulers with LRRL.
+Section 4.2 – LRRL with Learning Rate Schedulers.
   • Optimizer: Adam
@@
   • Initial learning rate of schedulers: $\eta_{0}=6.25\times 10^{-5}$
-Section 4.3– Comparing RMSProp and Adam Optimizers
+Section 4.3 – Continuous Action Spaces.
+  • Optimizer: Adam
+  • Bandit step-size: $\alpha=0.2$
+  • AdamRel/Initial learning rate: $3\times 10^{-4}$
+  • LRRL/CLR set of learning rates: $\mathcal{K}=\bigl\{1.2\times 10^{-4},1.8\times 10^{-4},2.5\times 10^{-4}\bigr\}$
+Section B.2 – Comparing adversarial and stochastic MAB algorithms.
+  • Optimizer: Adam
+  • MOSS/Exp3 set of learning rates: $\mathcal{K}=\bigl\{3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr\}$
+Section B.3– Comparing RMSProp and Adam Optimizers.
   • Optimizer: RMSProp for DQN baselines
   • Bandit step-size: $\alpha=0.2$
-  • LRR-DQN set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
-Section 4.4– LRRL with a Distributional RL Algorithm.
+  • LRRL-DQN set of learning rates: $\mathcal{K}=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr\}$
+Section B.4– Distributional RL: Implicit Quantile Networks (IQN).
   • Optimizer: Adam
@@
   • LRRL-IQN set of learning rates: $\mathcal{K}=\bigl\{2.5\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\bigr\}$
-Section B.2 – A Comparison between Multi-Armed Bandit Algorithms.
-  • Optimizer: Adam
-  • MOSS/Exp3 set of learning rates: $\mathcal{K}=\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
 Hyperparameters of the optimizers and deep RL algorithms.
@@
 The optimizers from the Optax library [13] are employed alongside the JAX [9] implementation of the DQN [28, 29] and IQN [12] algorithms, as provided by the Dopamine framework [10].
+For the MuJoCo tasks, we use the implementation provided by TorchRL [8].
 DQN and IQN hyperparameters Value
@@
 Environment steps per iteration (steps) 250 000
-Reward clipping [-1, 1]
 Network neurons per layer 32, 64, 64
@@
 $\beta_{2}$ decay 0.999
-Eps (DQN) 1.5$\times$10-4
-Eps (IQN) 3.125$\times$10-4
+[…]
+[…]
 RMSProp hyperparameters (DQN)
@@
 Centered False
-Eps 1.5$\times$10-4
-Table 1: Hyperparameters used in the experiments.
+[…]
+Table 2: Hyperparameters used in the experiments with discrete action spaces.
+PPO hyperparameters Value
+Discount factor ($\gamma$) 0.99
+Mini-batch size ($\mathcal{B}$) 64
+Epochs 10
+GAE 0.95
+Clip $\epsilon$ 0.2
+Critic coef. 0.25
+Entropy coef. 0.0
+Minimum environment steps ($\kappa$) 1 episode
+Network neurons per layer 64, 64
+Hardware (GPU) V100
+Table 3: Hyperparameters used in the experiments with continuous action spaces.